### v3

# Peer-induced Fairness: A Causal Approach for Algorithmic Fairness Auditing

Abstract
As artificial intelligence and automation increasingly permeate decision-making systems, ensuring algorithmic fairness has become critical. This paper addresses a fundamental question often overlooked: how to audit algorithmic fairness scientifically. It is crucial to discern whether adverse decisions stem from algorithmic discrimination or merely from the subjects’ insufficient capabilities. To tackle this, we develop an algorithmic fairness auditing framework, “peer-induced fairness”, that leverages counterfactual fairness and advanced causal inference techniques, such as the Single World Intervention Graph. Our approach transcends the typical trade-off between quantitative fairness measures and accuracy by aiming to assess algorithmic fairness at the individual level through peer comparisons and hypothesis testing, particularly in contexts like credit approval. This framework effectively addresses data scarcity and imbalance—frequent data quality challenges in traditional models, and is a uniquely model-agnostic and flexible self-audit tool for stakeholders and an external audit tool for regulators, in a plug-and-play fashion. Additionally, it offers explainable feedback for those who receive unfavourable decisions due to insufficient capabilities. We validate our framework in a practical context, highlighting the degree of algorithmic bias that arises, with 41.51% and 56.40% of subjects being either discriminated against or privileged respectively. It could also serve as a transparent, and adaptable tool suitable for diverse applications.
Key Words: Ethics in OR, Algorithmic Fairness, Causality, Counterfactual Fairness
1 Introduction
Algorithmic data-driven methods are extensively employed across a broad spectrum of fields, including healthcare, advertising, employment, supply chain, credit scoring, criminal justice (Kozodoi et al.,, 2022; Guan et al.,, 2020; Chen and Hooker,, 2022; Berk et al.,, 2017; Dwork et al.,, 2012; Lodi et al.,, 2023, 2024). These approaches are being adopted to replace human decision-making with the aim of reducing biases and ostensibly moving society towards greater equality. This is because algorithms and robots, as non-human entities, do not inherently possess the biases that can influence human judgement, thereby reducing the potential for discrimination. Admittedly, algorithms do not discriminate against groups of individuals; however, algorithmic data-driven models often rely on historical datasets that may contain significant biases. There are a number of potential areas of use of these datasets where algorithmic bias might arise. Therefore, to truly achieve the objective of reducing discrimination, it is crucial to develop algorithmic models that are specifically designed to handle and correct for biases within these datasets. Such models must effectively make decisions that are free from discrimination based on protected characteristics, such as gender or marital status (Lessmann et al.,, 2015; Kozodoi et al.,, 2022).
Such algorithmic fairness is not only applicable to people but exists in any context where an organisation may be treated unfairly in a decision process. As regulatory policies and documents, such as the General Data Protection Regulation (GDPR) (Voigt and Von Dem Bussche,, 2017) and the Markets in Financial Instruments Directive II (MiFID II) (Yeoh,, 2019), continue to evolve, algorithmic fairness auditing and monitoring becomes increasingly significant. The Europe Artificial Intelligence Act (Madiega,, 2021) (EU AI Act), the first comprehensive AI law in the world, requires that high-risk applications, such as credit scoring, to identify discrimination themselves. It also mandates that AI systems in these applications should be assessed before being put on the market and continuously throughout their life cycle. British Standards Institution (BSI) subsequently requires that AI providers ensure full compliance with the EU AI Act (British Standards Institution,, 2023). To support clients who will be regulated by this legislation, BSI emphasises the significance of algorithm auditing services, which are designed to help AI providers meet the necessary standards and regulatory requirements set by the EU AI Act (British Standards Institution,, 2023). Therefore, offering a reasonable tool for self-audit for stakeholders, as well as for external audits by regulators, is crucial.
However, a key challenge is obtaining precise and stable auditing results, especially given poor data quality, such as data scarcity and imbalance, which are common issues in many types of research. Additionally, it is crucial to distinguish whether the rejection of an individual or organisation is due to discrimination or inherent incapability.
In response, our paper employs a straightforward yet effective causal framework to audit the existence of algorithmic bias—that is, to determine whether certain individuals or organisations are being unfairly treated. The core concept involves comparing treatment in model outcomes among similarly situated individuals or organisations within the dataset. By analysing these comparisons, we aim to ascertain whether any individual or organisation is being unfairly treated by the algorithm.
Although the concept of comparing similar individuals is neither new nor complex, our framework, termed “peer-induced fairness” makes significant contributions to the field. First, this is the first framework to formalise a practical concept of “peer-induced fairness” specifically designed to audit algorithmic biases. Unlike traditional static measures of fairness, “peer-induced fairness” is an advanced framework that leverages counterfactual fairness (Wu et al.,, 2019) and causal inference techniques, such as Single World Intervention Graphs (SWIGs) (Richardson and Robins,, 2013) and peer observation theory (Li and Jain,, 2016; Ho and Su,, 2009). Stakeholders and regulators could use it as a bias audit tool for self-assessment and external assessment. Second, the core counterfactual comparison approach makes “peer-induced fairness” robust against data scarcity issues—a common challenge where protected groups are often underrepresented in datasets. This methodology circumvents the need for traditional statistical estimates within the protected group by utilising robust counterfactual statistics derived from well-represented peer groups. This approach allows us to use data from a single group, thereby making our method robust to population imbalance. Third, “peer-induced fairness” offers a transparent framework that could distinguish the subjects who are unfairly treated and merely not capable. It enables comparisons between an individual’s and their peers’ data, providing watch-out insights regarding the key features of why they are fairly treated but still rejected. Fourth, we validate our framework on access to finance for small and medium-sized enterprises (SMEs). Our application highlights in detail many advantages of using the framework which details unfairness and other aspects such as explainability. Given it relates to organisations, it expands the literature from predominantly individual-focused studies to include corporate entities, using firm size as a protected characteristic, given that smaller businesses are more likely to be denied loans and unfairly treated in accessing finance. Fifth, our criteria and framework are adaptable, offering potential for broader application across various fields and datasets. It is not only applicable to credit scoring for individuals but also extends to firms and can be adapted to any domain requiring fairness analysis. Its versatility stems from its ability to handle multifaceted scenarios independently of the specific fairness measures or the underlying problem because it assesses peers in a counterfactual manner.
The rest of our paper is structured as follows. Section 2 reviews the relevant literature. Section 3 starts from the counterfactual world and introduces the casual framework. Section 4 proposes the peer observation theory, corresponding peer identification process and “peer-induced fairness” framework. Section 5 and Section 6 demonstrate our experiment procedure and empirical results. Section 7 concludes.
2 Literature review
The theory and practice surrounding fairness have garnered increasing attention from both scholars and regulators (Federal Trade Commission,, 2023; Rohner,, 1979; Voigt and Von Dem Bussche,, 2017; Kehrenberg et al.,, 2020). The concept of algorithmic fairness in automated decision-making systems is notably complex and lacks a universally accepted definition. Several frameworks have been proposed to address this challenge (Dwork et al.,, 2012; Hardt et al.,, 2016).
Despite advancements in developing measures to uphold fairness criteria, there remains a considerable gap in the practical application of these frameworks. Traditionally, academic efforts have focused on transforming the concept of fairness into quantifiable definitions that address discrimination within specific datasets. However, those responsible for implementing these algorithms—such as practitioners, policymakers, and judicial figures—face significant challenges in choosing the most appropriate fairness definition to suit their unique circumstances (Kusner et al.,, 2017; Huang et al.,, 2020; Dixon et al.,, 2018; Foulds et al.,, 2020; Hickey et al.,, 2020). For example, the criteria for fairness required to address gender disparities may vary markedly from those needed for racial issues, and similarly, from the broader, non-demographic contexts, such as ensuring equitable treatment between large corporations and SMEs in credit approval processes (Lu and Calabrese,, 2023). It is impractical to adopt a single quantitative fairness definition as a universal solution for all sectors. Furthermore, it is essential to recognise that, despite their positive intentions, some fairness models can inadvertently increase discrimination (Kozodoi et al.,, 2022). This highlights the urgent need for a detailed and context-specific evaluation of fairness definitions, to ensure that the deployment of algorithmic decision-making systems genuinely contributes to reducing bias (Kusner et al.,, 2017).
In response to persistent issues in algorithmic decision-making, a causally-oriented approach to fairness has been advocated (Kusner et al.,, 2017), focusing on the relationships between protected features and data. Subsequent studies (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Chiappa,, 2019) have shown the efficacy of causal inference techniques in developing fair algorithms. However, counterfactual fairness encounters significant limitations, such as unidentifiability from observational data under certain conditions, which complicates the measurement of counterfactual outcomes (Wu et al.,, 2019). Additionally, the challenge of data scarcity often hinders decision-making processes intended to implement fairness constraints. Historical biases typically result in datasets where protected groups are underrepresented (Iosifidis and Ntoutsi,, 2018), thereby skewing the accuracy of fairness criteria and utility metrics. This imbalance, particularly the under-representation of minority groups in training data (i.e., representational disparity), leads to their diminished influence on model objectives (Hashimoto et al.,, 2018). As a result, biased measures of discrimination may emerge (Sha et al.,, 2023; Dablain et al.,, 2022). For example, in the finance sector, the availability of credit approval data for minority groups is substantially lower than for majority groups, complicating the fair assessment of creditworthiness—a critical aspect of many established fairness frameworks. Furthermore, the implementation of complex causal frameworks often depends on intricate causal graph assumptions and elaborate causal inferences, such as those detailed in (Chiappa,, 2019). These calculations are not only complex but also typically require Monte Carlo approximations. While these frameworks are adequate for group-level fairness analyses, they are less suited to addressing the needs of specific individuals or firms. Those seeking to enhance their chances of approval for future financing applications require more tangible explanations and actionable feedback than that which Monte Carlo approximations can provide. Regulatory authorities have consistently emphasised the necessity for transparent and explainable models to provide clear decision-making grounds (Chen et al.,, 2024; Voigt and Von Dem Bussche,, 2017). However, current explanation-related fairness criteria usually incorporate explainability into the fairness framework (Zhao et al.,, 2023; Hickey et al.,, 2020). There is still a gap in providing explanations of the fairness framework, which is crucial as it aids people in understanding the specific reasons behind the rejections. Therefore, this paper contributes to filling the gaps by designing a novel fairness framework using a causal lens, to stably audit algorithmic bias with group imbalance and data scarcity, and provide explanations to promote the transparency of our framework.
3 Counterfactual fairness and SWIGs
The attractiveness of counterfactual reasoning stems from its capacity to rigorously analyse causal relationships, unearth potential biases, and furnish methodologies for elucidating decisions made by models. Counterfactual reasoning critically examines and establishes causal connections by contemplating hypothetical scenarios under altered conditions (e.g., “If the individual were not a woman, would her application be approved for a loan?”). Counterfactual fairness is a concept that has been explored and represented in diverse forms within the academic literature (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Wu et al.,, 2019). In this paper, we adopt the general framework as described by Wu et al., (2019).
Let $S$ represent the set of protected features of an individual, which by definition, must not be subject to bias under any fairness doctrine. Additionally, let $\bm{Z}$ represent the set of unprotected features, with $\bm{X}\subseteq\bm{Z}$ specifying the subset of observable features for any given individual. The outcome of the decision-making process, potentially influenced by historical biases, is denoted by $Y$. We utilise a historical dataset $\mathcal{D}$, sampled from a distribution $\mathbb{P}(\bm{Z},S,Y)$, to train a classifier $f:(\bm{Z},S)\mapsto\hat{Y}$, where $\hat{Y}$ is the prediction generated by a machine learning algorithm aiming to estimate $Y$. The causal structure underlying the distribution $\mathbb{P}(\bm{Z},S,\hat{Y})$ is represented by a graph causal model $\mathcal{G}$.
Definition 1 (Counterfactual fairness).
Given a set of features $\bm{X}\subseteq\bm{Z}$, a classifier $f:(\bm{X},S)\mapsto\hat{Y}$ is counterfactually fair with respect to $\bm{X}$ if under any observable context $\bm{X}=\bm{x}$ and $S=s$,
$\mathbb{P}(\hat{Y}_{S\leftarrow s}=y|\bm{X}=\bm{x},S=s)=\mathbb{P}(\hat{Y}_{S\leftarrow s^{\prime}}=y|\bm{X}=\bm{x},S=s),$ (1)
for all $y$ and for any value $s^{\prime}$ attainable by $S$.
For a binary protected feature and a dichotomous decision outcome, a simplified version can be formulated.
Definition 2.
Given a set of features $\bm{X}\subseteq\bm{Z}$, a binary classifier $f:(\bm{X},S)\mapsto\hat{Y}$ is counterfactually fair with respect to $\bm{X}$ if under any observable context $\bm{X}=\bm{x}$ and $S=s_{-}$,
$\mathbb{P}(\hat{Y}_{S\leftarrow s_{-}}=1|\bm{X}=\bm{x},S=s_{-})=\mathbb{P}(\hat{Y}_{S\leftarrow s_{+}}=1|\bm{X}=\bm{x},S=s_{-}),$ (2)
for all $y$ and $S=\{s_{+},s_{-}\}$.
For illustrative purposes, imagine a scenario where individuals/organisations are evaluated for accessing finance using a predictive model, which determines the decision outcome, represented as $\hat{Y}$. Let us focus on a firm from the smallest SME (i.e., micro group), denoted by $s_{-}$ with a specific profile $\bm{x}$. The likelihood that this firm receives a favourable outcome is expressed as $\mathbb{P}(\hat{Y}|s_{-},\bm{x})$, which is equivalent to $\mathbb{P}(\hat{Y}_{S\leftarrow s_{-}}=1|S=s_{-},\bm{X}=\bm{x})$ by maintaining the firm’s protected feature (i.e., firms’ size) unaltered. Suppose, hypothetically, that this firm’s protected feature is changed from $s_{-}$ to $s_{+}$. The probability of a favourable outcome after such a counterfactual modification is denoted by $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$. Counterfactual fairness is achieved when the probabilities $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ and $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ are equal, suggesting that the treatment of the firm would remain consistent irrespective of their group membership. This condition underscores the essence of counterfactual fairness, where the decision-making process is indifferent to changes in the protected features of the firms.
$S$ $s_{-}$ $\bm{X}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{-}})$ $\bm{x}$ $Y({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{-}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{x}})$
(a) Actual Scenario: $\mathcal{G}(s_{-},\bm{x})$
$S$ $s_{+}$ $\bm{X}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{-}})$ $\bm{x}$ $Y({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{+}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{x}})$
(b) Counterfactual Scenario: $\mathcal{\tilde{G}}(s_{+},\bm{x})$
$S$ $s_{+}$ $\bm{X}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{+}})$ $\bm{x}^{\prime}$ $Y({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{+}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{x}^{\prime}})$
(c) Actual Scenario: $\mathcal{G}(s_{+},\bm{x}^{\prime})$
Figure 1: SWIGs for Graphical Causal Models (GCM). (a): The SWIG $\mathcal{G}(s_{-},\bm{x})$ represents the actual scenario for an individual with features $(s_{-},\bm{x})$. (b): The SWIG $\mathcal{\tilde{G}}(s_{+},\bm{x})$ illustrates the counterfactual scenario, assuming the individual’s protected feature changes from $s_{-}$ to $s_{+}$, while their other features $\bm{x}$ remain the same. (c): The SWIG $\mathcal{G}(s_{+},\bm{x}^{\prime})$ represents the actual scenario for an individual with features $(s_{+},\bm{x}^{\prime})$. The actual SWIG $\mathcal{G}(s_{-},\bm{x})$ corresponds to the conditional distribution $\hat{Y}_{s_{-}}|s_{-},\bm{x}$. Conversely, in the counterfactual SWIG $\mathcal{\tilde{G}}(s_{+},\bm{x})$ refers to $\hat{Y}_{s_{+}}|s_{-},\bm{x}$, denoting the outcome distribution had the individual been featured with $s_{+}$, given that the actual features are $(s_{-},\bm{x})$. Thus the directed link from $s_{+}$ to $\bm{X}(s_{-})$ is not the fact (shown in green colour). Note: $\mathcal{\tilde{G}}(s_{+},\bm{x})\neq\mathcal{G}(s_{+},\bm{x}^{\prime})$ because $\mathcal{\tilde{G}}(s_{+},\bm{x})$ is counterfactual scenario with actual features $(s_{-},\bm{x})$ and $\mathcal{G}(s_{+},\bm{x}^{\prime})$ is the fact with features $(s_{+},\bm{x}^{\prime})$.
A more nuanced comprehension of counterfactual fairness may be facilitated through the lens of SWIGs [^1] In SWIGs, black nodes represent random variables, while red nodes indicate fixed values, representing experimental interventions. Arrows depict causal relationships between variables. (Richardson and Robins,, 2013). Consider an individual belonging to a disadvantaged group $s_{-}$, characterised by features $\bm{x}$. The label $s_{-}$ could exert a direct influence on the outcome $Y$, or it may indirectly impact $Y$ through its effect on other observable features $\bm{X}$. If we postulate a counterfactual scenario in which the individual’s group designation changes from $s_{-}$ to $s_{+}$, the corresponding Graphical Causal Models (GCMs) for both actual and hypothetical situations can be depicted using SWIGs, as illustrated in Fig. 1. Counterfactual fairness is attained if the predictor, consistent with the actual GCM and the counterfactual GCM, yields identical probabilities for the outcome given the specific features $(s_{-},\bm{x})$.
Next, let us review some pivotal conclusions derived from the SWIGs as depicted in Panel (a) of Fig. 1 and propose some notations. A key aspect we will discuss is the factorisation properties of the joint distribution of all variables within a SWIG, applicable to any protected feature $s_{-},s_{+}$ and other features $\bm{x}$, which can be mathematically represented as follows:
$\mathcal{G}(s,\bm{x}):\mathbb{P}(S,\bm{X}(s),Y(s,\bm{x}))=\mathbb{P}(S)\cdot\mathbb{P}(\bm{X}(s))\cdot\mathbb{P}(Y(s,\bm{x})),s\in\{s_{-},s_{+}\}.$ (3)
Furthermore, the modularity property is observed where:
$\mathbb{P}(\bm{X}(s)=\bm{x})=\mathbb{P}(\bm{X}=\bm{x}|S=s),s\in\{s_{-},s_{+}\},$ (4)
$\mathbb{P}(Y(s,\bm{x})=y)=\mathbb{P}(Y=y|\bm{X}=\bm{x},S=s),s\in\{s_{-},s_{+}\},$ (5)
highlighting the left-hand side is the potential outcome while the right-hand side is the observational conditional probability. In the context of the counterfactual scenario with actual features $(s_{-},\bm{x})$ shown in Panel (b) of Fig. 1, a similar joint distribution is applicable:
$\mathcal{\tilde{G}}(s_{+},\bm{x}):\mathbb{P}(S,\bm{X}(s_{-}),Y(s_{+},\bm{x}))=\mathbb{P}(S)\cdot\mathbb{P}(\bm{X}(s_{-}))\cdot\mathbb{P}(Y(s_{+},\bm{x})).$ (6)
4 Peer-induced fairness with causal framework
While the concept of counterfactual fairness is theoretically straightforward and can be easily described, its application in practice is hampered by the challenges in identifying counterfactual outcomes from observational data in certain scenarios, as highlighted by Wu et al., (2019). Specifically, the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ as a potential outcome remains elusive for direct calculation due to its unidentifiability. To navigate this impediment and facilitate a feasible implementation of counterfactual fairness, we propose a practical approximation method that utilises peer comparison as an effective strategy.
4.1 Discrimination from peer comparisons
The phenomenon of discrimination, a ubiquitous aspect of daily life, is extensively explored within cognitive science literature. Research indicates that perceptions of discrimination are shaped not only by personal experiences but also through comparisons with peers who, despite possessing similar capabilities, skills, or knowledge, experience differential treatment, leading to missed opportunities. These perceptions are cultivated both through individual encounters and the lens of peer experiences. When an individual’s treatment aligns with that of their peer group, perceptions of being biased tend to diminish. Studies have shown that social and financial ties are more likely to form among individuals who share similarities in revenue levels, consumption behaviours, educational background, class, gender, race, or creditworthiness, illustrating a preference for homogeneity (Li et al.,, 2020; Haenlein,, 2011; Goel and Goldstein,, 2014; Wei et al.,, 2016).
4.2 Fairness through peer observations
Building on the concept of bias through peer comparisons discussed previously, we propose a more rigorous mathematical representation to demonstrate this idea effectively.
Consider an individual $A$ from a protected group with a protected status $S=s_{-}$ and other unprotected features $\bm{X}=\bm{x}$, denoted as $A=(s_{-},\bm{x})$. Assuming the protected and unprotected groups are comparable, if there exists a group of peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ from the unprotected group $S=s_{+}$, represented as $\{(s_{+},\bm{x}_{1}),(s_{+},\bm{x}_{2}),\cdots\}$, forming an $A$-oriented network. We use the expectation of the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})$ across these peers $\mathcal{C}$ to approximate the counterfactual $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$, mathematically expressed as
$\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})\approx\mathbb{E}_{(s_{+},\bm{x}_{j})\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})],$ (7)
where $\mathbb{E}{[\cdot]}$ is the expectation (or average) notation. This peer-based counterfactual approximation is intuitive, adhering to the non-discrimination principle where, ideally, the unobserved counterfactual probability aligns consistently with the average observed among peers. The method avoids the necessity for conventional statistical estimations within the protected group by employing resilient counterfactual statistics obtained from adequately represented peer groups. It adeptly addresses data scarcity within the protected group.
4.3 Peer definition and identification
Before initiating peer comparisons, we need to formulate the definition of peers.
Definition 3 ($\delta$-peer).
Let us consider an individual $A$ belonging to a protected group, characterised by a protected feature $S=s_{-}$ and a set of unprotected features $\bm{X}=\bm{x}_{0}$, represented as $A=(s_{-},\bm{x}_{0})$. Assuming there exists a set of individuals =$\mathcal{B}=\{B_{1},B_{2},\cdots\}$ from the unprotected group, where $B_{i}=(s_{+},\bm{x}_{i})$ for $i=1,2,\ldots$. An individual $C\in\mathcal{B}$ is defined as $\delta$-peer of $A$ if the difference in joint distributions between $C$’s actual SWIG, $\mathcal{G}(s_{+},\bm{x}_{j})$, and $A$’s counterfactual SWIG, $\mathcal{\tilde{G}}(s_{+},\bm{x}_{0})$, is less than a threshold $\delta$,
$\left|\mathbb{P}(\mathcal{G}(s_{+},\bm{x}_{j}))-\mathbb{P}(\mathcal{\tilde{G}}(s_{+},\bm{x}_{0}))\right|<\delta,$ (8)
where $\mathbb{P}(\mathcal{G}(s_{+},\bm{x}_{j}))=\mathbb{P}(S,\bm{X}(s_{+}),Y(s_{+},\bm{x}_{j}))$ and $\mathbb{P}(\mathcal{\tilde{G}}(s_{+},\bm{x}_{0}))=\mathbb{P}(S,\bm{X}(s_{-}),Y(s_{+},\bm{x}_{0}))$.
The concept of a peer in the graphical causal model is defined through the interrelations among three random variables: $S$, $\bm{X}$, and $Y$. For rigorous and unbiased comparisons, it is essential that a peer exhibits a joint distribution similar to the counterfactual scenario.
Despite the appealing theoretical foundation of the $\delta$-peer concept, its practical implementation encounters significant challenges. A primary obstacle is the difficulty in calculating $\mathbb{P}(Y(s_{+},\bm{x}))$ from Eq. (6) for the counterfactual case $(s_{+},\bm{x})$, which is crucial for assessing peer similarity in such contexts. This complication stems from the representation of $\bm{x}$ as the unprotected features for the protected group, where direct calculation of this probability is often unfeasible due to the absence of observational data. To address this and develop a more feasible approach for peer selection, we re-examine Eq. (3) and Eq. (6).
Since it is not feasible to directly derive $\mathcal{\tilde{G}}(s_{+},\bm{x})$ from observational data, we have no choice but use the information from $\mathcal{G}(s_{+},\bm{x})$ as a proxy for approximation, which has been discussed in Section 4.2. Upon comparing Eq. (3) and Eq. (6), the difference lies in the terms $\bm{X}$ and $Y$. Referring to Panel (a) of Fig. 1 and considering $\bm{x}_{0}$ as the observable unprotected features of an individual from the protected group $S=s_{-}$, we can compute $\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ using Bayes’ formula:
$\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ $=\mathbb{P}(\bm{X}=\bm{x}_{0}|S=s_{-})$
$=\frac{\mathbb{P}(\bm{X}=\bm{x}_{0})\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}_{0})}{\mathbb{P}(S=s_{-})}.$ (9)
Similarly, we can determine $\mathbb{P}(\bm{X}(s_{+})=\bm{x}_{0})$:
$\mathbb{P}(\bm{X}(s_{+})=\bm{x}_{0})$ $=\mathbb{P}(\bm{X}=\bm{x}_{0}|S=s_{+})$
$=\frac{\mathbb{P}(\bm{X}=\bm{x}_{0})\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})}{\mathbb{P}(S=s_{+})}.$ (10)
However, because $\bm{x}_{0}$ are the observable unprotected features for an individual from the protected group $S=s_{-}$, estimating $\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ directly is not feasible. Given that $S$ represents a binary set, we can infer:
$\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ $=1-\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}_{0}),$ (11)
$\mathbb{P}(S=s_{+})$ $=1-\mathbb{P}(S=s_{-}).$ (12)
We propose a unified notation for both $\mathbb{P}(\bm{X}(s_{-}))$ and $\mathbb{P}(\bm{X}(s_{+}))$ to streamline calculations and ensure consistency across analyses:
$\mathbb{P}(\bm{X}(s)=\bm{x})=\mathbb{P}(\bm{X}=\bm{x})\xi(s,\bm{x}),$ (13)
where $\xi(s,\bm{x})$ is defined as the identification coefficient (IC). This coefficient adjusts the probability values to reflect the conditions of being either a factual or counterfactual group, and is given by:
$\xi(s,\bm{x})=\begin{cases}\frac{1}{\mathbb{P}(S=s_{-})}\cdot\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}),&\text{if }s=s_{-},\\ \frac{1}{1-\mathbb{P}(S=s_{-})}\cdot(1-\mathbb{P}(S=s_{-}|\bm{X}=\bm{x})),&\text{if }s=s_{+}.\end{cases}$ (14)
Although direct evaluation of the joint distribution $\mathcal{\tilde{G}}(s_{+},\bm{x})$ is not feasible, we can facilitate the comparison by utilising the computable $\xi(s,\bm{x})$. This approach hinges on quantitative comparison and addresses the critical question: “How can peers be identified?”. Traditional methods often employ multi-dimensional matching to identify similar individuals within datasets, typically focusing on unprotected features $\bm{X}$. However, the causal impact of protected features $S$ on $\bm{X}$, coupled with the high dimensionality of $\bm{X}$, poses significant challenges to the efficacy of these conventional matching techniques. The complexity introduced by the curse of dimensionality makes the straightforward application of these methods problematic.
We propose a practical approach to implement a $\delta$-peer identification algorithm. The approach utilises information from the counterpart group, effectively addressing the issues of data scarcity and imbalance theoretically.
Theorem 1 ($\delta$-peer identification).
Consider an individual $A=(s_{-},\bm{x}_{0})$ and assuming there are a group of individuals =$\mathcal{B}=\{B_{1},B_{2},\cdots$ from unprotected group, where $B_{j}=(s_{+},\bm{x}_{j})$. An individual $C\in\mathcal{B}$ is identified as a $\delta$-peer of $A$ if:
$|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{j})|<\delta.$ (15)
Proof.
According to Definition 3, we have
$\quad\left|\mathbb{P}(\mathcal{G}(s_{+},\bm{x}_{j}))-\mathbb{P}(\mathcal{\tilde{G}}(s_{+},\bm{x}_{0}))\right|$
$=\Big{|}\mathbb{P}(S,\bm{X}(s_{+}),Y(s_{+},\bm{x}_{j}))-\mathbb{P}(S,\bm{X}(s_{-}),Y(s_{+},\bm{x}_{0}))\Big{|}$
$=\mathbb{P}(s_{+})\cdot\Big{|}\mathbb{P}(\bm{X}(s_{+}))\cdot\mathbb{P}(Y(s_{+},\bm{x}_{j}))-\mathbb{P}(\bm{X}(s_{-}))\mathbb{P}(Y(s_{+},\bm{x}_{0}))\Big{|}$
$=\mathbb{P}(s_{+})\cdot\Big{|}\mathbb{P}(\bm{x}_{j})\cdot\xi(s_{+},\bm{x}_{j})\cdot\mathbb{P}(Y|s_{+},\bm{x}_{j})-\mathbb{P}(\bm{x}_{0})\cdot\xi(s_{-},\bm{x}_{0})\cdot\mathbb{P}(Y|s_{+},\bm{x}_{0})\Big{|}$
$=\mathbb{P}(s_{+})\cdot\Big{|}\xi(s_{+},\bm{x}_{j})\cdot\mathbb{P}(Y,\bm{x}_{j}|s_{+})-\xi(s_{-},\bm{x}_{0})\cdot\mathbb{P}(Y,\bm{x}_{0}|s_{+})\Big{|}$
$=\mathbb{P}(s_{+})\cdot\mathbb{P}(Y,\bm{x}_{j}|s_{+})\cdot|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{j})|$
$\leq\mathbb{P}(s_{+})\cdot\mathbb{P}(Y,\bm{x}_{j}|s_{+})\cdot\delta$
$<\delta.$
The derivation of the second equation is underpinned by the factorisation property, as detailed in Eq. (3) and Eq. (6). The transition to the third equation leverages the modularity property, which is articulated in Eq. (5). The transition from $\mathbb{P}(\bm{X}(s_{+}))$ and $\mathbb{P}(\bm{X}(s_{-}))$ into $\mathbb{P}(\bm{x}_{j})\cdot\xi(s_{+},\bm{x})$ and $\mathbb{P}(\bm{x}_{0})\cdot\xi(s_{+},\bm{x}_{0})$ refer to Eq. (13). Regarding the fifth equation, it addresses the practical consideration of dealing with high-dimensional continuous variables in $\bm{X}$. Given the high-dimensional nature of $\bm{X}$, the probability of $\bm{X}$ equating to a specific value within this space is nominally small. Thus, for practical purposes, the distinction between $\mathbb{P}(Y,\bm{X}=\bm{x}_{j}|s_{+})$ and $\mathbb{P}(Y,\bm{X}=\bm{x}_{0}|s_{+})$ is considered negligible (i.e., $\mathbb{P}(Y,\bm{x}_{j}|s_{+})=\mathbb{P}(Y,\bm{x}_{0}|s_{+})$). Therefore, $C$ is considered as a peer of $A$ according to Definition 3. ∎
Consequently, we can generate an algorithm shown in Algorithm 1 to identify all peers in the dataset.
Input: A set of individuals $\{A\}=\{(s_{-},\bm{x}_{0})\}$ from the protected group, a set of individuals $\{B_{i}\}_{i=1}^{N}=\{(s_{+},\bm{x}_{i})\}$ from the unprotected group, a threshold $\delta$, and a minimum number of peers $U$.
Output: A subset of $\{B_{i}\}$ designated as $\delta$-peers of $A$, with each protected individual having at least $U$ peers.
1 foreach $A=(s_{-},\bm{x}_{0})$ do
2       Initialise an empty list of peers for $A$, denoted as $\text{Peers}_{A}$
3       Compute $\xi(s_{-},\bm{x}_{0})$ for $A$
4       foreach $B_{i}=(s_{+},\bm{x}_{i})$ in $\{B_{i}\}_{i=1}^{N}$ do
5             Compute $\xi(s_{+},\bm{x}_{i})$ for $B_{i}$
6             Calculate the difference $\Delta=|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{i})|$
7             if $\Delta<\delta$ then
8                   Add $B_{i}$ to $\text{Peers}_{A}$
9
10             end if
11
12       end foreach
13
14 end foreach
Algorithm 1 Identification of $\delta$-Peers for Protected Group Individuals
4.4 Peer-induced fairness
Following the idea of peer comparison, definition, and identification, we can now introduce the concept of peer-induced fairness.
Definition 4 ($(\delta,f)$-peer-induced fairness [^2] Although the term “peer-induced fairness” has been used in other contexts as noted by (Ho and Su,, 2009; Li and Jain,, 2016), our concept is novel in its reliance on a structured causal reasoning framework specifically tailored for classification tasks.).
Consider an individual $A=(s_{-},\bm{x}_{0})$ and assuming $A$ has a number of $\delta$-peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ where $C_{j}=(s_{+},\bm{x}_{j})$. $A$ is said to be fairly treated by the peers subject to $(\delta,f)$ if and only if
$\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})=\mathbb{E}_{C_{j}\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|C_{j})],$ (16)
where $\hat{Y}$ is the predictive outcome provided with the classifier $f$.
As discussed in previous sections, while we can directly estimate $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ from individual observations, estimating the expected value $\mathbb{E}_{C_{j}\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|C_{j})]$ presents challenges due to the limited number of observations available for $\delta$-peers. Consequently, we have to rely on observable peers to approximate the population mean. To formalise this, we introduce the random variable
$T_{j}=\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}).$ (17)
Upon examining the distribution of $T_{j}$, we find that it does not follow a normal distribution, with details presented in Supplementary Materials. Therefore, we randomly select a subset of peers and use the sample mean to estimate the population mean,
$\bar{T}=\frac{1}{K}\sum_{j=1}^{K}\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}),$ (18)
where $K$ is a large enough number of peers in the subset.
According to the Central Limit Theorem, the sample mean $\bar{T}$ follows a normal distribution, and thus $\mathbb{E}[\bar{T}]$, can be employed to estimate the overall predictive probabilities of favourable outcomes among peers, denoted as $\mathbb{E}[T]=\mu$. Based on this, we propose a proposition that a synthetic individual, defined using $IC$  [^3] Although the synthetic individual is defined by $IC$, the corresponding predictive favourable outcome probabilities calculation should follow Eq. (18)., can also be considered as a $\delta$-peer.
Proposition 1.
Let $A$ be an individual and $\mathcal{C}=\{C_{1},C_{2},\ldots\}$ denote all of $A$’s $\delta$-peers. Define a synthetic individual $\bar{T}_{i}$ using the average $IC$ of any subset $\mathcal{C}_{i}$ of $K$ peers, where $\mathcal{C}_{i}=\{C_{1}^{i},\ldots,C_{K}^{i}\}\subseteq\mathcal{C}$, $i\in\{1,2,\ldots,N\}$ and $C_{j}^{i}$ represents the $j$-th peer in the $i$-th selection with the unprotected feature $\bm{x}_{j}^{i}$. This synthetic individual $\bar{T}_{i}$ can also be considered as a $\delta$-peer of $A$.
Proof.
To demonstrate that the synthetic individual $\bar{T}_{i}$ qualifies as a $\delta$-peer of $A$, we compare $A$’s $IC$, $\xi(s_{-},\bm{x}_{0})$, against the average $IC$ of any $K$ peers of $A$, denoted as $\sum_{j=1}^{K}\xi(s_{+},\bm{x}_{j})/K$. The difference is calculated as follows:
$\quad\left|\xi(s_{-},\bm{x}_{0})-\frac{1}{K}\sum_{j=1}^{K}\xi(s_{+},\bm{x}_{j})\right|$
$=\frac{1}{K}\left|K\xi(s_{-},\bm{x}_{0})-\sum_{j=1}^{K}\xi(s_{+},\bm{x}_{j})\right|$
$=\frac{1}{K}\left|(\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{1}))+\cdots+(\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{K}))\right|$
$\leq\frac{1}{K}\sum_{j=1}^{K}\left|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{j})\right|$
$\leq\delta.$
This inequality shows that the average discrepancy between $A$’s $IC$ and that of $\bar{T}_{i}$ is within $\delta$. Hence, according to Theorem 1, $\bar{T}_{i}$ indeed qualifies as a $\delta$-peer of $A$. ∎
Consequently, by randomly selecting $K$ peers from the set of all observed $\delta$-peers $N$ times, we calculate the predictive favourable outcome probabilities $\bar{T}_{i}$ for each $i$-th selection using Eq. (18). Here
$\bar{T}_{i}=\frac{1}{K}\sum_{j=1}^{K}\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}^{i}).$ (19)
We then utilise the mean of the observed sample mean distribution, $\{\bar{T}_{i}\}_{i=1}^{N}$, which are all confirmed $\delta$-peers as per Proposition 1, to estimate the overall mean $\mu$ of favourable outcome probabilities among all peers.
4.5 Hypothesis testing for peer-induced fairness
Finally, to formalise the process of auditing whether an individual in a protected group is subjected to algorithmic bias, we propose a hypothesis-testing framework. This framework is predicated on an appropriate threshold for peer identification $\delta$ and a specific classifier $f$. It aims to test whether the sample mean distribution $\{\bar{T}_{i}\}_{i=1}^{N}$ is statistically equivalent to $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$. Since $\bar{T}_{i}$ follows a normal distribution and $N$ is a large enough number, our hypothesis is consistent with the standard $z$-test, which is designed to evaluate the presence of algorithmic bias statistically.
  • $H_{0}$ (Null Hypothesis): The individual $A=(s_{-},\bm{x}_{0})$ is equally treated according to $(\delta,f)$-“peer-induced fairness” criterion,
    $H_{0}:\mathbb{E}[\bar{T}_{i}]=\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0}).$ (20)
  • $H_{1}$ (Alternative Hypothesis): The individual $A$ is subject to algorithmic bias under $(\delta,f)$-“peer-induced fairness” criterion, which is evidenced by a significant disparity in treatment compared to their unprotected peers,
    $H_{1}:\mathbb{E}[\bar{T}_{i}]\neq\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0}).$ (21)
Furthermore, it is also feasible to consider two additional scenarios: checking whether the individual is algorithmically discriminated against, where $H_{2}:\mathbb{E}[\bar{T}_{i}]<\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$, or algorithmically benefited, where $H_{3}:\mathbb{E}[\bar{T}_{i}]>\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$.
5 Experiment setup
Fairness concepts apply not only to decision-making concerning groups of people but also to broader domains, such as companies within the economic system. The banking loan approval algorithm that uses historical data often places micro-firms at a disadvantage due to their smaller size and limited historical records compared to larger firms (Cenni et al.,, 2015). Consequently, these smaller entities may face higher interest rates or be more likely to be denied loans, despite their growth potential.
Therefore, to illustrate the usefulness of our novel “peer-induced fairness” framework, we consider a real example using SMEs data. The collected data is from the UK Archive Small and Medium-Sized Enterprise Finance Monitor (BDRC Continental,, 2023). The dataset compiles survey information on SMEs  [^4] SMEs included in this survey meet the four criteria: 1) employ no more than 250 individuals, 2) have an annual turnover not exceeding £25 million, 3) do not operate as social enterprises or non-profit organisations, and 4) are not owned by another company by more than 50%. spanning from 2011Q1 to 2023Q4, with approximately 4,500 telephone interviews conducted per quarter across the UK. Each interview provides insights into the experiences of SMEs with external financing over the past 12 months, including their anticipated future financial needs and perceived obstacles to growth. It also details the characteristics of the SMEs and their owners or managers. To avoid redundancy, we selected survey results from 2012Q4 to 2020Q2. We focused on 15 important features identified from the literature (Sun et al.,, 2021; Calabrese et al.,, 2022; Cowling et al.,, 2016, 2022, 2012) (refer to Table 1 for details). These features demonstrate various dimensions of the loan application process. After filtering out data points with more than 20% missing features, we obtained a dataset comprising 4,159 data points for our analysis. The details of data cleaning are presented in Supplementary Materials.
Table 1: Description of features.
Feature Category Value Percentage
risk multivariate & ordinal minimal 19.59%
low 43.11%
average 25.98%
above average 11.31%
principal multivariate & nominal construction 6.64%
agriculture, hunting and forestry 10.82%
fishing 12.01%
health and social work 12.62%
hotels and restaurants 11.68%
manufacturing 8.69%
real estate, renting and business activities 16.68%
transport, storage and communication 9.63%
wholesale/retail 11.23%
other community, social and personal service 9.63%
legal status multivariate & nominal sole proprietorship 4.88%
partnership 10.57%
limited liability partnership 7.50%
limited liability company 77.05%
loss or profit multivariate & ordinal loss 86.07%
broken even 8.69%
profit 5.25%
turnover growth rate multivariate & ordinal grown more than 20% 13.69%
grown but by less than 20% 40.33%
stayed the same 33.69%
declined 12.30%
funds injection binary no 67.17%
yes 32.83%
credit purchase binary no 18.48%
yes 81.52%
startups binary no 97.5%
yes 2.5%
previous turn-down binary no 90.94%
yes 9.06%
London & South East binary yes 76.39%
no 23.61%
business innovation binary no 40.16%
yes 59.84%
product/service development binary no 70.25%
yes 29.75%
regular management account binary no 19.06%
yes 80.94%
written plan binary no 37.58%
yes 62.42%
finance qualification binary no 45.66%
yes 54.34%
For this analysis, we designate the observed features listed in Table 1 as $X$. We treat the firm size as the protected attribute $S$, defined by a combination of the number of employees and annual turnover. This categorisation results in 1,719 micro-firms ($s=s_{-}$) classified as protected and 2,440 non-micro firms ($s=s_{+}$) as unprotected [^5] Micro-firms are defined as those with fewer than 10 employees and an annual turnover of less than £2 million (Sun et al.,, 2021).. We consider the outcome of bank loan applications, denoted as $Y$, due to the significant role of bank loans in SME financing (Sun et al.,, 2021). The dataset records 3,391 approvals ($y=1$) and 768 rejections ($y=0$), highlighting the decisions faced by SMEs in securing financial support.
In subsequent analyses, we focus on the 1,719 micro-firms to determine if they have experienced algorithmic bias using our “peer-induced fairness” framework. Initially, it is essential to identify each firm’s peers by computing the $IC$ as specified in Algorithm 1. Without loss of generality, we set the default $\delta$ to 0.3 times the standard deviation of the micro-firms $ICs$. This flexible threshold can be adjusted based on the dataset of the specific research field. The specific robustness tests are presented in Supplementary Materials. However, direct estimation of $IC$ from observed data is challenging, necessitating the use of a fitted model. For simplicity, we employ a logistic classifier to estimate the probability of an individual being labelled as part of the protected group, with performance and robustness test in Supplementary Materials.
Following Proposition 1, we then utilise the expectation of the observed sample mean distribution to estimate the overall mean of all peers. As a standard approach, we randomly sample $N=100$ times, and each sample consists of $K=30$ data points. We consider only those micro-firms with more than $U=35$ peers; firms with fewer than 35 peers are labelled as “Unknown”. Again, direct estimation of the mean from Eq. (18) is not feasible, requiring the use of a predictive classification model. By default, we use a logistic classifier with performance in Supplementary Materials, although any classification model could be substituted. The robustness of the classification model is demonstrated in Supplementary Materials. The data are typically split into training (80%) and testing (20%) sets, with hyper-parameters optimised via grid search and 5-fold cross-validation. The model yielding the highest AUC value is selected for predictions on the target $Y$.
Ultimately, we conduct hypothesis tests (i.e., $H_{0}$, $H_{2}$, $H_{3}$) to compare the mean approval likelihood of the peers against that of the micro-firms, thereby identifying potential algorithmic bias, discrimination and privilege. The significance level for the hypothesis test is set at $\alpha=5\%$. Due to the presence of “Unknown” micro-firms, hypothesis testing is performed on a subset of the data, consisting of 1007 data points.
6 Experiment results
In this section, we present the experimental results derived from the SMEs data to demonstrate the efficacy of our “peer-induced fairness” framework. These results illustrate how the framework operates in practice and underscore its potential for auditing bias in algorithmic decision-making.
6.1 Algorithmic fairness auditing
With the evolution of algorithmic fairness methods and the increasing regulatory demands for data protection and transparency in decision-making processes, there is a growing emphasis on applying advanced methodologies to ensure fairness in decision systems. For instance, in the banking sector, it is becoming common to self-scrutinise or externally audit decision-making processes to determine whether they meet established fairness criteria or if they continue to exhibit significant algorithmic bias.
Following the steps outlined in Section 5, we have identified algorithmic bias within this SMEs dataset. The scatter plot (refer to Fig. 2) comparing micro-firms to their peers regarding approval likelihood reveals that only 2.48% of them are treated fairly, indicating significant disparities in the credit approval system. 97.52% of them experience algorithmic bias, with 41.51% of micro-firms experiencing discrimination. An intriguing observation is that the remaining 56.40% of micro-firms, despite being under-represented, benefit from the decision system, receiving approval likelihoods higher than the average for their peers.
Figure 2: Comparative analysis of loan approval likelihood for micro-firms against peers. The black dashed 45-degree line, denoting $Y=X$, symbolises perfect fairness. Red and orange data points represent micro-firms with approval likelihoods significantly lower or higher, respectively than the average of their peers. Blue points denote no significant difference.
To identify the specific extent of discrimination and privilege faced by each micro-firm, we compare the approval likelihood difference between a given micro-firm $A=(s-,\bm{x_{0}})$ and its peers. For micro-firms with a higher likelihood of approval, we allow for greater tolerance when assessing extreme algorithmic bias, adjusting the standard based on each firm’s approval likelihood. Specifically, we consider a micro-firm to experience extreme algorithmic bias if the absolute difference exceeds 0.1 times its own approval likelihood. Mathematically, this is expressed as $|\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})-\mathbb{E}[\bar{T}_{i}]|>0.1\times\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$. A negative difference indicates discrimination, while a positive difference signifies privilege. This approach ensures the flexibility of the standard, making it suitable for firms in different situations. Specifically, 26.71% of micro-firms experience substantial discrimination, with their approval likelihood markedly lower than that of their peers, as shown in Panel (a)-(c) in Fig. 3, at both group and individual levels. 32.17% of micro-firms are extremely privileged, as shown in Panel (g)-(i). Even though algorithmic privilege might seem beneficial for micro-firms, neither scenario is desirable. We advocate for transparency and fairness in decision-making processes. Arbitrary or opaque factors influencing decisions are contrary to the principles of fairness and should be rigorously addressed to ensure equitable treatment across all applicants.
Figure 3: Comparative analysis of loan approval likelihood for micro-firms under each algorithmic treatment category against peers. (a)-(c): Extremely discriminated (ED) micro-firms. (d)-(f): Fairly treated (FT) micro-firms. (g)-(i): Extremely privileged (EP) micro-firms. The second column provides the approval likelihood comparison between these micro-firms and their peers at the group level. The third column provides the comparison, at the individual level, between the selected micro-firm in the first column and its peers.
It is important to emphasise that our framework is a tool for audits by regulators and stakeholders, aiming to detect algorithmic bias. In the credit loan application, rejected customers are particularly concerned about whether they were treated fairly, while regulators and banks require detailed results to audit the fairness of their models. Therefore, our framework also includes detailed information on accepted applicants. Additionally, without compromising generalisability to other research areas, it is crucial to focus on all applicants.
We also validate our framework by investigating the connection between accessing finance outcomes and disparities in algorithmic bias. Among these markedly discriminated micro-firms, 52.42% were denied loans, whereas only 9.97% of their peers faced rejection, highlighting a significant disparity in rejection rates. The rejection rate of micro firms decreases and that of their peers increases with the diminished discrimination. The difference in rejection rates between micro-firms and their peers also decreases. The rejection rates of peers fluctuate around the rejection rate of fairly treated micro-firms. This fluctuation indicates that within the category, some micro-firms experience higher rejection rates compared to their peers, while others experience lower rejection rates, illustrating a gradual convergence in rejection rates across categories with less pronounced discrimination. Notably, even the lowest peer rejection rate surpasses that of micro-firms in the extremely privileged category, where micro-firms experience the lowest rejection rates, as in Fig. 4. These findings, derived from our bias audit based on financing outcomes prediction, align with the observed financing results. This congruence further validates the utility of our framework in accurately reflecting disparities and biases in the loan approval process. Further details of the degree of algorithmic bias are provided in Supplementary Materials.
Figure 4: Rejection rates of micro-firms across algorithmic treatment categories and their peers. The algorithmic treatment categories include extremely discriminated (ED), fairly treated (FT), and extremely privileged (EP). Each category includes multiple micro-firms with a single rejection rate, shown as histograms, while the rejection rate of peers of each micro-firm in this category is represented in the black line with error bars to indicate variability.
From the analysis presented, it is evident that our “peer-induced fairness” framework not only identifies disparities in algorithmic fairness but also facilitates the visual representation of individual-level discrepancies across all users in the dataset. This capability allows for clear visualisation of algorithmic fairness, where discrimination or benefit is readily distinguishable. Such insights are invaluable not only for regulatory purposes but also for verifying the effectiveness of algorithmic fairness models. Furthermore, we subjected all results to robustness tests, varying the level of peer identification threshold, model fitting selection, and prediction algorithms to ensure the integrity of our findings.
6.2 Data scarcity and imbalance
Data scarcity and imbalance significantly influence the performance of advanced machine learning models due to the potential for inaccurate parameter estimation (Chen et al.,, 2024; Lessmann et al.,, 2015). This issue is particularly pronounced in the field of algorithmic fairness, where the representation of minority groups is often limited compared to majority groups. This discrepancy caused by the poor data quality, subsequently affects the assessment of algorithmic fairness.
Our “peer-induced fairness” framework addresses these challenges uniquely. Unlike traditional models that rely heavily on the data from the protected group, our framework bases all parameter estimations on peers identified within the unprotected group. This group typically possesses ample data points, effectively mitigating issues related to data scarcity and group imbalance, making our framework robust theoretically.
We investigate the robustness of our peer-induced fairness framework by evaluating the percentages of unfairly treated ($PUT$) protected individuals or organisations and the invariant outcome ratio ($IOR$) under varying levels of imbalance. The imbalance ratio, $\omega$, is defined as the proportion of samples in the protected class:
$\omega=\frac{\#(S=s_{-})}{\#(S=s_{+})+\#(S=s_{-})},$
where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality of a set. A perfectly balanced dataset corresponds to $\omega=50\%$. The $PUT$ is calculated as the number of unfairly treated individuals or organisations divided by the total number of selected subjects in the experiments with different $\omega$. The $IOR$ is computed as the number of selected individuals or organisations in the experiment with $\omega$ that have unchanged predictive outcomes compared to the original experiment ($\omega=41.33\%$) divided by the number of commonly selected subjects in both the experiment with $\omega$ and the original experiment.
In SMEs experiment, building upon the default settings outlined in Section 5, we explore the influence of varying imbalance ratios by randomly selecting subsets of the original dataset with controlled imbalance ratios. Specifically, we evaluate performance at imbalance ratios of $\omega=\{36.33\%,31.33\%,26.33\%,21.33\%,16.33\%,11.33\%\}$. By decreasing the percentage of micro-firms in these subsets, we assess the framework’s performance across different levels of imbalance. To mitigate the effects of randomness inherent in subset selection, the process is repeated five times. The detailed procedure is presented in Supplementary Materials.
The results are visualised in Fig. 5 and demonstrate the robustness of our framework. From the view of $PUT$, the small error bars across all the imbalance levels suggest the results across the five repetitions are highly consistent. This observation underscores the robustness of our “peer-induced framework” to imbalanced datasets. From the view of $IOR$, it is approximately 95% and remains stable across different imbalance levels. This aligns with our expectations, as the framework does not rely on data from the minority group but rather leverages information from the unprotected group, leading to inherent robustness. The small error bars suggest that for imbalance ratios greater than or equal to 16.33%, the results regarding $IOR$ are also highly consistent.
Figure 5: Percentage of unfairly treated micro-firms and invariant outcome ratio at different group imbalance levels. The imbalance level is represented on the x-axis as a percentage, ranging from 11.33% to 36.33%. The left y-axis shows the percentage of unfairly treated micro-firms (blue line), while the right y-axis displays the invariant outcome ratio (red line) as the imbalance level changes from the initial level to other levels.
These findings underscore the stability of our “peer-induced fairness” framework, distinguishing itself from others by effectively addressing the data scarcity and imbalance issues. Given the widespread nature of these issues, our framework holds considerable significance for researchers investigating algorithmic fairness and data imbalance. An alternative computation method is also provided in Supplementary Materials to ensure robustness.
6.3 Explainable fairness discovery
In our previous experiments, we were able to distinctly classify individuals from the protected group into two categories: fairly-treated and unfairly-treated groups. Our analysis now turns to those who were rejected while still fairly treated, to understand the reasons behind their rejections by comparing their features with those of their peers. Additionally, the “peer-induced fairness” framework allows us to provide a clear watch-out list of a series of features.
Given the existence of accepted peers as the counterfactual instances with positive accessing finance outcomes, the micro-firm which is fairly treated should originally have the same outcomes. Our framework identifies the feature differences between each rejected while fairly treated micro-firm and its accepted peers by hypothesis testing, with details presented in Supplementary Materials. For each feature, we summarise the percentage of these micro-firms that perform significantly worse than their accepted peers. We consider some actionable and key features to identify and understand these discrepancies, as in Fig. 6. The descriptions for each feature value are shown in Supplementary Materials. Results show that even though none of them have been rejected previously and only 25% of them perform worse on financial qualifications and written plans, banks generally prioritise the financial and business health of firms. 75% of these micro-firms invest excessively in business innovation and have lower risk ratings. Besides, half of them invest in product/service development and have lower profits. The uncertain returns and high risks associated with innovation lead to the failure or commercial non-viability of most innovative products (Coad and Rao,, 2008; Hall,, 2002; Freel,, 2007), exacerbating already poor-performing risk indicators. The worse performance on these key features makes banks cautious about the long-term financial sustainability of these firms. It also reflects the capability of these micro-firms, negatively affecting their loan approvals.
Figure 6: Comparative analysis of key features for rejected while fairly treated micro-firms vs. accepted peers. The x-axis represents the selected key attributes being analysed, including finance qualification (FQ), written plan (WP), previous turn-down (PT), loss or profit (LP), risk (RI), product/service development (PS), business innovation (BI). The y-axis represents the percentage of those micro-firms with significantly worse performance than their accepted peers on each feature.
This exploration identifies the differences between micro-firms and their peers for each feature and summarises the percentage of micro-firms that perform worse on each feature. This explainable analysis not only enhances the transparency of our framework but also supports regulators and stakeholders in understanding the specific challenges most incapable micro-firms face and highlights the features that they need to watch out for and pay extra attention to.
7 Conclusion
In this paper, we introduce a novel fairness framework within a causal framework, termed “peer-induced fairness”, as a bias auditing tool for internal and external assessment, in a plug-and-play fashion. It applies the principles of counterfactual fairness, stipulating that the average treatment of an individual should align with that of their peers. We identify peers based on similar joint distributions but resort to $IC$ due to the unidentifiability of the counterfactual distribution. The framework requires equal treatment with peers. This approach effectively tackles data scarcity and group imbalance by utilising robust counterfactual statistics derived from well-represented peer groups, thereby ensuring more stable bias auditing. Besides, based on the essence of peer comparison, we could also provide an explainable watch-out list for those who receive unfavourable decisions due to insufficient capabilities, promoting transparency of our method. We have applied this framework to SMEs, but it has the potential for a generalisation to other research domains.
By experimenting on SMEs data, our research findings reveal that by comparing micro-firms with their peers, banks and regulators can effectively audit algorithmic bias. Specifically, only 2.48% of micro-firms are treated fairly. 41.51% and 56.40% of micro-firms are either discriminated against or privileged respectively. Even though some micro-firms benefit from algorithmic favouritism, it is essential to ensure equitable treatment across all applicants. Nearly half of the micro-firms experiencing extreme discrimination are rejected, with a rejection rate, compared to only 9.97% among their peers. This difference diminishes and becomes negative as discrimination lessens and shifts towards privilege. Up to 95% of micro-firms maintained consistent auditing results despite changing imbalance levels, demonstrating the stability of our framework with data scarcity and imbalance issues. Additionally, the approach highlights the key features that financial institutions need to pay more attention to and rejected micro-firms may need to address, whilst clearly fairly treating micro-firms. The comparison could distinguish the bias and incapability faced by micro-firms, helping banks and regulators understand the specific issues these firms encounter.
Our research is significant for researchers who aim to scientifically audit algorithmic bias. This tool could perform with common data quality issues, like data scarcity and imbalance, ensuring accuracy and stability in measuring unfairness. Besides, our framework could also distinguish those incapable protected individuals from being biased individuals, preventing improving the treatment of less capable individuals at the expense of the treatment of capable, unprotected individuals. Besides, our empirical analysis is based on SMEs. Unlike previous studies that mainly focused on individual loans, our research extends the focus to the firm level. The protected feature is the firms’ size, differing from the traditional focus on person-level characteristics. This approach broadens the perspective of fairness. Despite the focus on SMEs loan approval, the fairness audit framework proposed can be applicable to other domains where algorithmic unfairness may occur, especially those suffering from group imbalance and data scarcity.
In the domain of fairness research, class imbalance is also a crucial issue involving poor data quality. It refers to the imbalance on the target label, leading the model to favour the majority class, thereby affecting the overall performance of the model and the fairness measure. Previous studies discussing the impact of class imbalance on fairness focus on the education domain (Sha et al.,, 2022, 2023), with the issue remaining unexplored in the credit scoring domain. This is particularly important because different datasets exhibit significant variations in features, labels, missing values, and sample sizes. Besides, Iosifidis et al., (2021, 2019) have proposed fair models for addressing class imbalance, but there is still a lack of a fairness framework that explicitly considers class imbalance naturally. Exploring the impact of class imbalance on algorithmic fairness measures, and developing fairness criteria related to class imbalance, are crucial for all domains reliant on precise data-driven decision-making.
Acknowledgments
The authors of this manuscript would like to thank Prof.Raffaella Calabrese and Dr.Yizhe Dong for their assistance and support in the discussion and research direction.

### v4

@@
 Abstract
-As artificial intelligence and automation increasingly permeate decision-making systems, ensuring algorithmic fairness has become critical.
-This paper addresses a fundamental question often overlooked: how to audit algorithmic fairness scientifically.
-It is crucial to discern whether adverse decisions stem from algorithmic discrimination or merely from the subjects’ insufficient capabilities.
-To tackle this, we develop an algorithmic fairness auditing framework, “peer-induced fairness”, that leverages counterfactual fairness and advanced causal inference techniques, such as the Single World Intervention Graph.
-Our approach transcends the typical trade-off between quantitative fairness measures and accuracy by aiming to assess algorithmic fairness at the individual level through peer comparisons and hypothesis testing, particularly in contexts like credit approval.
-This framework effectively addresses data scarcity and imbalance—frequent data quality challenges in traditional models, and is a uniquely model-agnostic and flexible self-audit tool for stakeholders and an external audit tool for regulators, in a plug-and-play fashion.
-Additionally, it offers explainable feedback for those who receive unfavourable decisions due to insufficient capabilities.
-We validate our framework in a practical context, highlighting the degree of algorithmic bias that arises, with 41.51% and 56.40% of subjects being either discriminated against or privileged respectively.
-It could also serve as a transparent, and adaptable tool suitable for diverse applications.
-Key Words: Ethics in OR, Algorithmic Fairness, Causality, Counterfactual Fairness
+With the European Union’s Artificial Intelligence Act taking effect on 1 August 2024, high-risk AI applications must adhere to stringent transparency and fairness standards.
+This paper addresses a crucial question: how can we scientifically audit algorithmic fairness?
+Current methods typically remain at the basic detection stage of auditing, without accounting for more complex scenarios.
+We propose a novel framework, “peer-induced fairness”, which combines the strengths of counterfactual fairness and peer comparison strategy, creating a reliable and robust tool for auditing algorithmic fairness.
+Our framework is universal, adaptable to various domains, and capable of handling different levels of data quality, including skewed distributions.
+Moreover, it can distinguish whether adverse decisions result from algorithmic discrimination or inherent limitations of the subjects, thereby enhancing transparency.
+This framework can serve as both a self-assessment tool for AI developers and an external assessment tool for auditors to ensure compliance with the EU AI Act.
+We demonstrate its utility in small and medium-sized enterprises’ access to finance, uncovering significant unfairness—41.51% of micro-firms face discrimination compared to non-micro firms.
+These findings highlight the framework’s potential for broader applications in ensuring equitable AI-driven decision-making.
+Key Words: Algorithmic fairness, AI auditing, Causality, Counterfactual fairness, Small and medium-sized enterprises, Credit scoring
 1 Introduction
-Algorithmic data-driven methods are extensively employed across a broad spectrum of fields, including healthcare, advertising, employment, supply chain, credit scoring, criminal justice (Kozodoi et al.,, 2022; Guan et al.,, 2020; Chen and Hooker,, 2022; Berk et al.,, 2017; Dwork et al.,, 2012; Lodi et al.,, 2023, 2024).
-These approaches are being adopted to replace human decision-making with the aim of reducing biases and ostensibly moving society towards greater equality.
-This is because algorithms and robots, as non-human entities, do not inherently possess the biases that can influence human judgement, thereby reducing the potential for discrimination.
-Admittedly, algorithms do not discriminate against groups of individuals; however, algorithmic data-driven models often rely on historical datasets that may contain significant biases.
-There are a number of potential areas of use of these datasets where algorithmic bias might arise.
-Therefore, to truly achieve the objective of reducing discrimination, it is crucial to develop algorithmic models that are specifically designed to handle and correct for biases within these datasets.
-Such models must effectively make decisions that are free from discrimination based on protected characteristics, such as gender or marital status (Lessmann et al.,, 2015; Kozodoi et al.,, 2022).
-Such algorithmic fairness is not only applicable to people but exists in any context where an organisation may be treated unfairly in a decision process.
-As regulatory policies and documents, such as the General Data Protection Regulation (GDPR) (Voigt and Von Dem Bussche,, 2017) and the Markets in Financial Instruments Directive II (MiFID II) (Yeoh,, 2019), continue to evolve, algorithmic fairness auditing and monitoring becomes increasingly significant.
-The Europe Artificial Intelligence Act (Madiega,, 2021) (EU AI Act), the first comprehensive AI law in the world, requires that high-risk applications, such as credit scoring, to identify discrimination themselves.
-It also mandates that AI systems in these applications should be assessed before being put on the market and continuously throughout their life cycle.
-British Standards Institution (BSI) subsequently requires that AI providers ensure full compliance with the EU AI Act (British Standards Institution,, 2023).
-To support clients who will be regulated by this legislation, BSI emphasises the significance of algorithm auditing services, which are designed to help AI providers meet the necessary standards and regulatory requirements set by the EU AI Act (British Standards Institution,, 2023).
-Therefore, offering a reasonable tool for self-audit for stakeholders, as well as for external audits by regulators, is crucial.
-However, a key challenge is obtaining precise and stable auditing results, especially given poor data quality, such as data scarcity and imbalance, which are common issues in many types of research.
-Additionally, it is crucial to distinguish whether the rejection of an individual or organisation is due to discrimination or inherent incapability.
-In response, our paper employs a straightforward yet effective causal framework to audit the existence of algorithmic bias—that is, to determine whether certain individuals or organisations are being unfairly treated.
-The core concept involves comparing treatment in model outcomes among similarly situated individuals or organisations within the dataset.
-By analysing these comparisons, we aim to ascertain whether any individual or organisation is being unfairly treated by the algorithm.
-Although the concept of comparing similar individuals is neither new nor complex, our framework, termed “peer-induced fairness” makes significant contributions to the field.
-First, this is the first framework to formalise a practical concept of “peer-induced fairness” specifically designed to audit algorithmic biases.
-Unlike traditional static measures of fairness, “peer-induced fairness” is an advanced framework that leverages counterfactual fairness (Wu et al.,, 2019) and causal inference techniques, such as Single World Intervention Graphs (SWIGs) (Richardson and Robins,, 2013) and peer observation theory (Li and Jain,, 2016; Ho and Su,, 2009).
-Stakeholders and regulators could use it as a bias audit tool for self-assessment and external assessment.
-Second, the core counterfactual comparison approach makes “peer-induced fairness” robust against data scarcity issues—a common challenge where protected groups are often underrepresented in datasets.
-This methodology circumvents the need for traditional statistical estimates within the protected group by utilising robust counterfactual statistics derived from well-represented peer groups.
-This approach allows us to use data from a single group, thereby making our method robust to population imbalance.
-Third, “peer-induced fairness” offers a transparent framework that could distinguish the subjects who are unfairly treated and merely not capable.
-It enables comparisons between an individual’s and their peers’ data, providing watch-out insights regarding the key features of why they are fairly treated but still rejected.
-Fourth, we validate our framework on access to finance for small and medium-sized enterprises (SMEs).
-Our application highlights in detail many advantages of using the framework which details unfairness and other aspects such as explainability.
-Given it relates to organisations, it expands the literature from predominantly individual-focused studies to include corporate entities, using firm size as a protected characteristic, given that smaller businesses are more likely to be denied loans and unfairly treated in accessing finance.
-Fifth, our criteria and framework are adaptable, offering potential for broader application across various fields and datasets.
-It is not only applicable to credit scoring for individuals but also extends to firms and can be adapted to any domain requiring fairness analysis.
-Its versatility stems from its ability to handle multifaceted scenarios independently of the specific fairness measures or the underlying problem because it assesses peers in a counterfactual manner.
-The rest of our paper is structured as follows.
-Section 2 reviews the relevant literature.
-Section 3 starts from the counterfactual world and introduces the casual framework.
-Section 4 proposes the peer observation theory, corresponding peer identification process and “peer-induced fairness” framework.
-Section 5 and Section 6 demonstrate our experiment procedure and empirical results.
-Section 7 concludes.
-2 Literature review
-The theory and practice surrounding fairness have garnered increasing attention from both scholars and regulators (Federal Trade Commission,, 2023; Rohner,, 1979; Voigt and Von Dem Bussche,, 2017; Kehrenberg et al.,, 2020).
-The concept of algorithmic fairness in automated decision-making systems is notably complex and lacks a universally accepted definition.
-Several frameworks have been proposed to address this challenge (Dwork et al.,, 2012; Hardt et al.,, 2016).
-Despite advancements in developing measures to uphold fairness criteria, there remains a considerable gap in the practical application of these frameworks.
-Traditionally, academic efforts have focused on transforming the concept of fairness into quantifiable definitions that address discrimination within specific datasets.
-However, those responsible for implementing these algorithms—such as practitioners, policymakers, and judicial figures—face significant challenges in choosing the most appropriate fairness definition to suit their unique circumstances (Kusner et al.,, 2017; Huang et al.,, 2020; Dixon et al.,, 2018; Foulds et al.,, 2020; Hickey et al.,, 2020).
-For example, the criteria for fairness required to address gender disparities may vary markedly from those needed for racial issues, and similarly, from the broader, non-demographic contexts, such as ensuring equitable treatment between large corporations and SMEs in credit approval processes (Lu and Calabrese,, 2023).
-It is impractical to adopt a single quantitative fairness definition as a universal solution for all sectors.
-Furthermore, it is essential to recognise that, despite their positive intentions, some fairness models can inadvertently increase discrimination (Kozodoi et al.,, 2022).
-This highlights the urgent need for a detailed and context-specific evaluation of fairness definitions, to ensure that the deployment of algorithmic decision-making systems genuinely contributes to reducing bias (Kusner et al.,, 2017).
-In response to persistent issues in algorithmic decision-making, a causally-oriented approach to fairness has been advocated (Kusner et al.,, 2017), focusing on the relationships between protected features and data.
-Subsequent studies (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Chiappa,, 2019) have shown the efficacy of causal inference techniques in developing fair algorithms.
-However, counterfactual fairness encounters significant limitations, such as unidentifiability from observational data under certain conditions, which complicates the measurement of counterfactual outcomes (Wu et al.,, 2019).
-Additionally, the challenge of data scarcity often hinders decision-making processes intended to implement fairness constraints.
-Historical biases typically result in datasets where protected groups are underrepresented (Iosifidis and Ntoutsi,, 2018), thereby skewing the accuracy of fairness criteria and utility metrics.
-This imbalance, particularly the under-representation of minority groups in training data (i.e., representational disparity), leads to their diminished influence on model objectives (Hashimoto et al.,, 2018).
-As a result, biased measures of discrimination may emerge (Sha et al.,, 2023; Dablain et al.,, 2022).
-For example, in the finance sector, the availability of credit approval data for minority groups is substantially lower than for majority groups, complicating the fair assessment of creditworthiness—a critical aspect of many established fairness frameworks.
-Furthermore, the implementation of complex causal frameworks often depends on intricate causal graph assumptions and elaborate causal inferences, such as those detailed in (Chiappa,, 2019).
-These calculations are not only complex but also typically require Monte Carlo approximations.
-While these frameworks are adequate for group-level fairness analyses, they are less suited to addressing the needs of specific individuals or firms.
-Those seeking to enhance their chances of approval for future financing applications require more tangible explanations and actionable feedback than that which Monte Carlo approximations can provide.
-Regulatory authorities have consistently emphasised the necessity for transparent and explainable models to provide clear decision-making grounds (Chen et al.,, 2024; Voigt and Von Dem Bussche,, 2017).
-However, current explanation-related fairness criteria usually incorporate explainability into the fairness framework (Zhao et al.,, 2023; Hickey et al.,, 2020).
-There is still a gap in providing explanations of the fairness framework, which is crucial as it aids people in understanding the specific reasons behind the rejections.
-Therefore, this paper contributes to filling the gaps by designing a novel fairness framework using a causal lens, to stably audit algorithmic bias with group imbalance and data scarcity, and provide explanations to promote the transparency of our framework.
-3 Counterfactual fairness and SWIGs
-The attractiveness of counterfactual reasoning stems from its capacity to rigorously analyse causal relationships, unearth potential biases, and furnish methodologies for elucidating decisions made by models.
-Counterfactual reasoning critically examines and establishes causal connections by contemplating hypothetical scenarios under altered conditions (e.g., “If the individual were not a woman, would her application be approved for a loan?”).
-Counterfactual fairness is a concept that has been explored and represented in diverse forms within the academic literature (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Wu et al.,, 2019).
-In this paper, we adopt the general framework as described by Wu et al., (2019).
+The increasing adoption of data-driven methodologies across various sectors has heightened global concerns about algorithmic bias.
+These concerns are underscored by the recent enactment of the European Union’s Artificial Intelligence Act (EU AI Act), effective from 1 August 2024 (Madiega,, 2024).
+The Act represents the world’s first comprehensive legal framework for Artificial Intelligence (AI).
+It imposes stringent requirements on high-risk AI applications, such as credit scoring systems, mandating the rigorous identification and mitigation of discrimination risks.
+Additionally, the Act requires that these AI systems undergo thorough assessments both prior to deployment and continuously throughout their operational life cycle, thereby ensuring sustained compliance with transparency and fairness standards.
+The implementation of the EU AI Act necessitates the urgent development of a universal and transparent tool capable of ensuring compliance with rigorous standards.
+This task is critical not only for regulators but for everyone involved in the AI community.
+A key challenge lies in scientifically assessing the fairness of decisions made by AI models—a question that remains at the forefront of ethical AI development.
+Over the past decade, the importance of algorithmic fairness has increasingly gained recognition as a vital component of responsible technology use and ethical AI development.
+Despite this growing awareness, a substantial body of literature highlights the trade-offs between accuracy and fairness (Kim et al.,, 2023; Huang et al.,, 2020; Guldogan et al.,, 2022; Dixon et al.,, 2018; Foulds et al.,, 2020; Chen et al.,, 2022; Hickey et al.,, 2020; Dwork et al.,, 2012; Hardt et al.,, 2016).
+However, there remains a significant gap in the development of robust auditing frameworks specifically designed to assess and monitor algorithmic bias.
+Some studies have proposed approaches for auditing algorithmic fairness with data-driven models (Cherian and Candès,, 2024; Brundage et al.,, 2020; Xue et al.,, 2020; Tramer et al.,, 2017; Si et al.,, 2021; Yan and Zhang,, 2022).
+However, these approaches primarily focus on fairness detection based on internally accessible data and models.
+When regulators obtain data externally, the underlying algorithmic models and decision-making processes are often unknown, necessitating a reassessment of whether the decisions produced by AI developers’ models are fair.
+Besides, although identifying the presence of bias is necessary, it represents only the initial step in the comprehensive auditing of AI systems, and it is insufficient for ensuring robust and reliable auditing frameworks.
+The two key factors concerning the complexities inherent in the real world—universality and transparency—are often overlooked.
+With respect to universality, a mature auditing framework should be adaptable to datasets with different characteristics.
+For example, a critical challenge in the development of auditing tools is the issue of data quality, particularly data scarcity and imbalance, which are pervasive in real-world datasets (Lessmann et al.,, 2015; Chen et al.,, 2024; Sha et al.,, 2023; Dablain et al.,, 2022).
+Historical biases frequently result in the under-representation of protected groups within datasets (Iosifidis and Ntoutsi,, 2018).
+This imbalance, particularly the under-representation of minority groups in training data (i.e., representational disparity), leads to their diminished influence on model objectives and reduces their influence on model objectives (Hashimoto et al.,, 2018).
+Consequently, biased discrimination measures and unreliable audit outcomes may emerge (Sha et al.,, 2023; Dablain et al.,, 2022; Sha et al.,, 2022; Yan et al.,, 2020).
+A particularly concerning issue arises when an individual or organisation is flagged as discriminated against in one dataset but deemed privileged in another due to varying degrees of imbalance.
+Such inconsistencies pose significant challenges to conducting universal audits.
+Current frameworks frequently assume that data is balanced across different groups, an assumption rarely met in practice.
+Some studies propose oversampling techniques, such as SMOTE, to address data imbalance (Sha et al.,, 2023, 2022; Yan et al.,, 2020).
+However, these methods can inadvertently “smooth out” critical features within the data, thereby altering the intricate relationships between variables (Chen et al.,, 2024), which may ultimately lead to distorted auditing outcomes.
+Such flawed outcomes can be disastrous for regulators, companies, and third-party auditors, as they could introduce even greater risks.
+Audits based on unrealistic assumptions may fail to detect genuine instances of algorithmic discrimination, allowing unfair practices to persist.
+Furthermore, inaccurate audits could lead to misguided adjustments in algorithms, potentially worsening performance or introducing new biases, thus compounding the original issues.
+If stakeholders, including the public, perceive the auditing process as flawed or unreliable, it can erode confidence in both the regulators and the entities being audited.
+In terms of transparency, both regulatory authorities and researchers have consistently underscored the importance of transparent and explainable models that provide clear justifications for decision-making (Chen et al.,, 2024; Voigt and Von Dem Bussche,, 2017).
+Current fairness-related criteria often incorporate explainability into the fairness assessment (Zhao et al.,, 2023; Hickey et al.,, 2020; Chen et al.,, 2022).
+However, a gap remains in delivering clear and understandable explanations.
+This gap is critical, as understanding the reasons behind rejections is essential for ensuring that decisions are perceived as fair.
+An opaque auditing framework can similarly lead both the audited entities and the public to perceive the audit process as incredible, thereby undermining its credibility and public confidence.
+Therefore, an effective auditing framework must elucidate the reasons for adverse decisions, particularly to distinguish whether such decisions are due to discrimination or inherent limitations.
+We propose a reliable and robust audit framework that solves the above issues, embodying both universality and transparency.
+It is a causally-oriented approach to fairness (Kusner et al.,, 2017).
+Compared to traditional static fairness criteria, a causally-oriented approach offers a more effective means of addressing real-world challenges.
+This approach is particularly valuable for practitioners, policymakers, and judicial authorities tasked with implementing algorithms designed to mitigate discrimination.
+Selecting an appropriate fairness definition that aligns with the specific nuances of each scenario can be a significant challenge.
+For instance, the parameters for fairness in addressing gender disparities may differ substantially from those relevant to racial issues, or from considerations in broader, non-demographic contexts, such as ensuring equitable treatment between large corporations and SMEs in credit approval processes.
+Thus, applying a single quantitative fairness definition as a universal remedy across all sectors is impractical.
+This reasoning underpins our reliance on the causal framework, previous studies (Gastwirth,, 1997; Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Chiappa,, 2019; Imai et al.,, 2023, 2013) have demonstrated the efficacy of causal inference techniques, with counterfactual causal inference standing out as particularly prominent.
+Counterfactual reasoning critically examines and establishes causal connections by contemplating hypothetical scenarios under altered conditions (e.g., “If the applicant were not a female, would the application be approved for a loan?”).
+While counterfactual approaches to fairness have been previously suggested (Kusner et al.,, 2017), a key limitation of counterfactual fairness is the unidentifiability from observational data (Wu et al.,, 2019).
+To address this, our framework proves that individuals or organisations with similar joint distributions could be identified as counterfactual instances.
+We identify these counterfactual instances as peers.
+We are motivated by the peer comparison perception (Li and Jain,, 2016), which involves the differential treatment experienced by individuals compared to their peers.
+When an individual’s treatment is consistent with that of their peer group, perceptions of bias tend to diminish.
+Building on this insight, we introduce a novel concept termed “peer-induced fairness”, which leverages the strengths of counterfactual fairness while addressing its limitations, thereby creating a more reliable and robust tool for auditing algorithmic fairness.
+This framework also serves as a valuable self-assessment tool, which is increasingly crucial for the AI community in the development of products that must comply with the EU AI Act.
+Beyond its general contributions to the field of algorithmic fairness auditing and self-assessment, our paper has the following particular contributions: First, to the best of our knowledge, our framework is the first to formalise a practical concept of “peer-induced fairness” specifically designed to audit algorithmic biases.
+This comprehensive framework goes beyond the initial stage of basic detection, enabling users to evaluate externally obtained data without accessing the decision-making process or the underlying algorithmic model.
+Second, our framework is universal and versatile in handling different levels of data quality, including datasets with highly skewed distributions of protected attributes issues often overlooked or inadequately addressed in previous studies.
+Third, when it is considered as a self-assessment tool, the “peer-induced fairness” framework not only provides conclusions from self-assessment but also offers insightful explanations through peer comparisons, enhancing transparency and explainability.
+Fourth, we demonstrate the practical utility of the “peer-induced fairness” framework in uncovering algorithmic fairness issues related to small and medium-sized enterprises (SMEs) access to finance, particularly in scenarios where AI systems are used for decision-making.
+To the best of our knowledge, this is the first study of algorithmic bias on SMEs’ access to finance [^1] Lu and Calabrese, (2023) proposed a method for assessing the discrimination in ground truth $Y$ in SMEs’ access to finance, rather than in the algorithmic predictions ($\hat{Y}$) made by AI systems.
+It also highlights the effectiveness of our framework as a self-assessment tool in real-world applications.
+Besides, our framework is universal, applicable across different domains and capable of addressing various types of protected attributes.
+For example, it expands the protected attribute from the customer level (e.g., gender and race) to the organisation level (i.e., firm size).
+The remainder of this paper is organised as follows.
+Section 2 begins with an overview of the foundational concepts of counterfactual fairness and causal framework, including the representation of Single World Intervention Graphs (SWIGs).
+In Section 3, we introduce our “peer-induced fairness” framework in a step-by-step manner.
+Sections 4 and 5 detail our experimental procedures and present the empirical results, using the example of SMEs’ access to finance in the UK.
+Finally, concluding remarks and further discussions are provided in Section 6.
+2 Counterfactual fairness and its representation
+Before presenting our framework, it is essential to introduce some key concepts related to counterfactual fairness and its representation.
+Various forms of counterfactual fairness have been proposed in the academic literature (Pfohl et al.,, 2019; Kim et al.,, 2021; Kusner et al.,, 2017; Wu et al.,, 2019).
+In this paper, we adopt the general framework outlined by Wu et al., (2019).
 Let $S$ represent the set of protected features of an individual, which by definition, must not be subject to bias under any fairness doctrine.
@@
 for all $y$ and $S=\{s_{+},s_{-}\}$.
-For illustrative purposes, imagine a scenario where individuals/organisations are evaluated for accessing finance using a predictive model, which determines the decision outcome, represented as $\hat{Y}$.
-Let us focus on a firm from the smallest SME (i.e., micro group), denoted by $s_{-}$ with a specific profile $\bm{x}$.
-The likelihood that this firm receives a favourable outcome is expressed as $\mathbb{P}(\hat{Y}|s_{-},\bm{x})$, which is equivalent to $\mathbb{P}(\hat{Y}_{S\leftarrow s_{-}}=1|S=s_{-},\bm{X}=\bm{x})$ by maintaining the firm’s protected feature (i.e., firms’ size) unaltered.
-Suppose, hypothetically, that this firm’s protected feature is changed from $s_{-}$ to $s_{+}$.
+For illustrative purposes, imagine a scenario - loan applications using a predictive model, which determines the decision outcome, represented as $\hat{Y}$.
+Let us focus on an application by a female, denoted by $s_{-}$ with a specific profile $\bm{x}$.
+The likelihood that this applicant received a favourable outcome is expressed as $\mathbb{P}(\hat{Y}|s_{-},\bm{x})$, which is equivalent to $\mathbb{P}(\hat{Y}_{S\leftarrow s_{-}}=1|S=s_{-},\bm{X}=\bm{x})$ by maintaining the protected feature (i.e., gender) unaltered.
+Suppose, hypothetically, that this applicant’s protected feature is changed from $s_{-}$ to $s_{+}$.
 The probability of a favourable outcome after such a counterfactual modification is denoted by $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$.
-Counterfactual fairness is achieved when the probabilities $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ and $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ are equal, suggesting that the treatment of the firm would remain consistent irrespective of their group membership.
-This condition underscores the essence of counterfactual fairness, where the decision-making process is indifferent to changes in the protected features of the firms.
+Counterfactual fairness is achieved when the probabilities $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ and $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ are equal, suggesting that the treatment of the applicant would remain consistent irrespective of their membership.
+This condition underscores the essence of counterfactual fairness, where the decision-making process is indifferent to changes in the protected features.
+A more nuanced comprehension of counterfactual fairness may be facilitated through the lens of SWIGs (Richardson and Robins,, 2013).
+Consider an individual belonging to a disadvantaged group $s_{-}$, characterised by features $\bm{x}$.
+The label $s_{-}$ could exert a direct influence on the outcome $Y$, or it may indirectly impact $Y$ through its effect on other observable features $\bm{X}$.
+If we postulate a counterfactual scenario in which the individual’s group designation changes from $s_{-}$ to $s_{+}$, the corresponding Graphical Causal Models (GCMs) for both actual and hypothetical situations can be depicted using SWIGs, as illustrated in Figure 1.
+Counterfactual fairness is attained if the predictor, consistent with the actual GCM and the counterfactual GCM, yields identical probabilities for the outcome given the specific features $(s_{-},\bm{x})$.
 $S$ $s_{-}$ $\bm{X}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{-}})$ $\bm{x}$ $Y({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s_{-}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{x}})$
@@
 Figure 1: SWIGs for Graphical Causal Models (GCM).
+The nodes with black border represent random variables, while red ones indicate fixed values of random variables, representing experimental interventions.
+Arrows depict causal relationships between variables.
 (a): The SWIG $\mathcal{G}(s_{-},\bm{x})$ represents the actual scenario for an individual with features $(s_{-},\bm{x})$.
@@
 Note: $\mathcal{\tilde{G}}(s_{+},\bm{x})\neq\mathcal{G}(s_{+},\bm{x}^{\prime})$ because $\mathcal{\tilde{G}}(s_{+},\bm{x})$ is counterfactual scenario with actual features $(s_{-},\bm{x})$ and $\mathcal{G}(s_{+},\bm{x}^{\prime})$ is the fact with features $(s_{+},\bm{x}^{\prime})$.
-A more nuanced comprehension of counterfactual fairness may be facilitated through the lens of SWIGs [^1] In SWIGs, black nodes represent random variables, while red nodes indicate fixed values, representing experimental interventions.
-Arrows depict causal relationships between variables.
-(Richardson and Robins,, 2013).
-Consider an individual belonging to a disadvantaged group $s_{-}$, characterised by features $\bm{x}$.
-The label $s_{-}$ could exert a direct influence on the outcome $Y$, or it may indirectly impact $Y$ through its effect on other observable features $\bm{X}$.
-If we postulate a counterfactual scenario in which the individual’s group designation changes from $s_{-}$ to $s_{+}$, the corresponding Graphical Causal Models (GCMs) for both actual and hypothetical situations can be depicted using SWIGs, as illustrated in Fig.
-1.
-Counterfactual fairness is attained if the predictor, consistent with the actual GCM and the counterfactual GCM, yields identical probabilities for the outcome given the specific features $(s_{-},\bm{x})$.
-Next, let us review some pivotal conclusions derived from the SWIGs as depicted in Panel (a) of Fig.
-1 and propose some notations.
+Next, let us review some pivotal conclusions derived from the SWIGs as depicted in Figure 1 (a) and propose some notations.
 A key aspect we will discuss is the factorisation properties of the joint distribution of all variables within a SWIG, applicable to any protected feature $s_{-},s_{+}$ and other features $\bm{x}$, which can be mathematically represented as follows:
@@
 highlighting the left-hand side is the potential outcome while the right-hand side is the observational conditional probability.
-In the context of the counterfactual scenario with actual features $(s_{-},\bm{x})$ shown in Panel (b) of Fig.
-1, a similar joint distribution is applicable:
+In the context of the counterfactual scenario with actual features $(s_{-},\bm{x})$ shown in Figure 1 (b), a similar joint distribution is applicable:
 $\mathcal{\tilde{G}}(s_{+},\bm{x}):\mathbb{P}(S,\bm{X}(s_{-}),Y(s_{+},\bm{x}))=\mathbb{P}(S)\cdot\mathbb{P}(\bm{X}(s_{-}))\cdot\mathbb{P}(Y(s_{+},\bm{x})).$ (6)
-4 Peer-induced fairness with causal framework
-While the concept of counterfactual fairness is theoretically straightforward and can be easily described, its application in practice is hampered by the challenges in identifying counterfactual outcomes from observational data in certain scenarios, as highlighted by Wu et al., (2019).
+While the above concept of counterfactual fairness is theoretically straightforward and can be easily described, its application in practice is hampered by the challenges in identifying counterfactual outcomes from observational data in certain scenarios, as highlighted by Wu et al., (2019).
 Specifically, the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$ as a potential outcome remains elusive for direct calculation due to its unidentifiability.
-To navigate this impediment and facilitate a feasible implementation of counterfactual fairness, we propose a practical approximation method that utilises peer comparison as an effective strategy.
+This creates a significant challenge for regulators, making it difficult to implement algorithmic bias auditing.
+3 Peer-induced fairness with causal method
+In this section, we propose a practical approximation method that utilises peer comparison as an effective strategy, in order to navigate the above impediment and facilitate a feasible implementation of counterfactual fairness.
+We then introduce our “peer-induced fairness” framework and algorithmic bias-detecting methods after providing the peer identification concept.
 3.1 Discrimination from peer comparisons
@@
 Research indicates that perceptions of discrimination are shaped not only by personal experiences but also through comparisons with peers who, despite possessing similar capabilities, skills, or knowledge, experience differential treatment, leading to missed opportunities.
-These perceptions are cultivated both through individual encounters and the lens of peer experiences.
+These perceptions are cultivated both through individual encounters and the lens of peer experiences (Li and Jain,, 2016).
 When an individual’s treatment aligns with that of their peer group, perceptions of being biased tend to diminish.
 Studies have shown that social and financial ties are more likely to form among individuals who share similarities in revenue levels, consumption behaviours, educational background, class, gender, race, or creditworthiness, illustrating a preference for homogeneity (Li et al.,, 2020; Haenlein,, 2011; Goel and Goldstein,, 2014; Wei et al.,, 2016).
+The insights from cognitive science highlight the importance of understanding how comparative experiences shape perceptions of discrimination among individuals and organisations.
+These groups may not only perceive but also actually experience biased outcomes when compared to their peers.
+Such perceived or real biases could erode trust in automated systems and, more broadly, undermine confidence in the regulators intended to ensure fairness.
+To mitigate these risks, regulators need to ensure that auditing frameworks are sensitive to these subtler forms of discrimination, which can arise from differences in treatment relative to peers.
 3.2 Fairness through peer observations
@@
 Assuming the protected and unprotected groups are comparable, if there exists a group of peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ from the unprotected group $S=s_{+}$, represented as $\{(s_{+},\bm{x}_{1}),(s_{+},\bm{x}_{2}),\cdots\}$, forming an $A$-oriented network.
-We use the expectation of the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})$ across these peers $\mathcal{C}$ to approximate the counterfactual $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$, mathematically expressed as
+We use the expectation of the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})$ across these peers $\mathcal{C}$ to approximate the counterfactual $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$, mathematically expressed as:
 $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})\approx\mathbb{E}_{(s_{+},\bm{x}_{j})\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})],$ (7)
@@
 3.3 Peer definition and identification
-Before initiating peer comparisons, we need to formulate the definition of peers.
+While the counterfactual predictive probability is provided in Eq.
+(7), a key question remains: “What is a peer” in mathematical terms?
+A precise definition and identification of peers are crucial before initiating peer comparisons, as they ensure accurate and reliable assessments.
+This clarity is essential for regulators to effectively audit potential biases in algorithmic decision-making.
 Definition 3 ($\delta$-peer).
-Let us consider an individual $A$ belonging to a protected group, characterised by a protected feature $S=s_{-}$ and a set of unprotected features $\bm{X}=\bm{x}_{0}$, represented as $A=(s_{-},\bm{x}_{0})$.
+Let an individual $A$ belong to a protected group, characterised by a protected feature $S=s_{-}$ and a set of unprotected features $\bm{X}=\bm{x}_{0}$, represented as $A=(s_{-},\bm{x}_{0})$.
 Assuming there exists a set of individuals =$\mathcal{B}=\{B_{1},B_{2},\cdots\}$ from the unprotected group, where $B_{i}=(s_{+},\bm{x}_{i})$ for $i=1,2,\ldots$.
@@
 where $\mathbb{P}(\mathcal{G}(s_{+},\bm{x}_{j}))=\mathbb{P}(S,\bm{X}(s_{+}),Y(s_{+},\bm{x}_{j}))$ and $\mathbb{P}(\mathcal{\tilde{G}}(s_{+},\bm{x}_{0}))=\mathbb{P}(S,\bm{X}(s_{-}),Y(s_{+},\bm{x}_{0}))$.
-The concept of a peer in the graphical causal model is defined through the interrelations among three random variables: $S$, $\bm{X}$, and $Y$.
-For rigorous and unbiased comparisons, it is essential that a peer exhibits a joint distribution similar to the counterfactual scenario.
-Despite the appealing theoretical foundation of the $\delta$-peer concept, its practical implementation encounters significant challenges.
-A primary obstacle is the difficulty in calculating $\mathbb{P}(Y(s_{+},\bm{x}))$ from Eq.
-(6) for the counterfactual case $(s_{+},\bm{x})$, which is crucial for assessing peer similarity in such contexts.
-This complication stems from the representation of $\bm{x}$ as the unprotected features for the protected group, where direct calculation of this probability is often unfeasible due to the absence of observational data.
+In a graphical causal model, the concept of a peer is defined through the interrelations among three random variables: $S$, $\bm{X}$, and $Y$.
+To enable rigorous and unbiased comparisons, it is crucial that a peer exhibits a joint distribution similar to the counterfactual scenario.
+However, since the counterfactual scenario is inherently unobservable, it is rarely possible to observe the exact same $\bm{X}$ and $Y$ in another group defined by different protected attributes.
+Although we have the mathematical representation of a $\delta$-peer in Definition 3, its practical implementation faces significant challenges.
+A major obstacle is the difficulty in calculating $\mathbb{P}(Y(s_{+},\bm{x}))$ from Eq.
+(6) for the counterfactual scenario $(s_{+},\bm{x})$, which is essential for evaluating peer similarity.
+This difficulty arises because $\bm{x}$ represents the unprotected features for the protected group, where the direct calculation of this probability is often infeasible due to the lack of observational data.
 To address this and develop a more feasible approach for peer selection, we re-examine Eq.
@@
 (6), the difference lies in the terms $\bm{X}$ and $Y$.
-Referring to Panel (a) of Fig.
-1 and considering $\bm{x}_{0}$ as the observable unprotected features of an individual from the protected group $S=s_{-}$, we can compute $\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ using Bayes’ formula:
+Referring to Figure 1 (a) and considering $\bm{x}_{0}$ as the observable unprotected features of an individual from the protected group $S=s_{-}$, we can compute $\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ using Bayes’ formula:
 $\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ $=\mathbb{P}(\bm{X}=\bm{x}_{0}|S=s_{-})$
@@
 However, because $\bm{x}_{0}$ are the observable unprotected features for an individual from the protected group $S=s_{-}$, estimating $\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ directly is not feasible.
-Given that $S$ represents a binary set, we can infer:
-$\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ $=1-\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}_{0}),$ (11)
-$\mathbb{P}(S=s_{+})$ $=1-\mathbb{P}(S=s_{-}).$ (12)
-We propose a unified notation for both $\mathbb{P}(\bm{X}(s_{-}))$ and $\mathbb{P}(\bm{X}(s_{+}))$ to streamline calculations and ensure consistency across analyses:
+Given that $S$ represents a binary set, we can infer $\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})=1-\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}_{0})$.
+We can rewrite $\mathbb{P}(\bm{X}(s_{-}))$ and $\mathbb{P}(\bm{X}(s_{+}))$ in a unified representation,
 $\mathbb{P}(\bm{X}(s)=\bm{x})=\mathbb{P}(\bm{X}=\bm{x})\xi(s,\bm{x}),$ (11)
@@
 $|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{j})|<\delta.$ (13)
+Theorem 1 provides a sufficient condition for Definition 3, with the proof detailed in Appendix A.
+Based on this, we propose using the IC for peer identification as a practical alternative to the infeasible joint distribution.
+By enhancing the identification of suitable peers, regulators can effectively audit potential biases in decision-making systems in real-world scenarios using feasible methods.
+More practically, we can also implement the idea as an algorithm shown in Appendix B to identify all peers in the dataset step by step.
+This algorithm, by applying a similarity threshold $\delta$, is grounded in cognitive science perception of discrimination, ensuring that peers are selected for meaningful comparison based on their IC similarities to the protected individual.
+3.4 Peer-induced fairness
+Following the idea of peer comparison, definition, and identification, we can now introduce the concept of “peer-induced fairness”.
+Definition 4 ($(\delta,f)$-peer-induced fairness [^2] Although the term “peer-induced fairness” has been used in other contexts, as noted by (Ho and Su,, 2009; Li and Jain,, 2016), our work is distinct.).
+Consider an individual $A=(s_{-},\bm{x}_{0})$ and assuming $A$ has a number of $\delta$-peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ where $C_{j}=(s_{+},\bm{x}_{j})$.
+$A$ is said to be fairly treated by the peers subject to $(\delta,f)$ if and only if
+$\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})=\mathbb{E}_{C_{j}\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|C_{j})],$ (14)
+where $\hat{Y}$ is the predictive outcome provided with the classifier $f$.
+As discussed in previous sections, while we can directly estimate $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ from individual observations, estimating the expected value $\mathbb{E}_{C_{j}\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|C_{j})]$ presents challenges due to the limited number of observations available for $\delta$-peers.
+Consequently, we have to rely on observable peers to approximate the population mean.
+To formalise this, we introduce the random variable
+$T_{j}=\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}).$ (15)
+Upon examining the distribution of $T_{j}$, we find that it does not always follow a normal distribution, with details presented in Supplementary Materials.
+Therefore, we randomly select a subset of peers and use the sample mean to estimate the population mean,
+$\bar{T}=\frac{1}{K}\sum_{j=1}^{K}T_{j},$ (16)
+where $K$ is a large enough number of peers in the subset.
+According to the Central Limit Theorem, the sample mean $\bar{T}$ follows a normal distribution, and thus $\mathbb{E}[\bar{T}]$, can be employed to estimate the overall predictive probabilities of favourable outcomes among peers, denoted as $\mathbb{E}[T]=\mu$.
+Based on this, we propose a proposition that a synthetic individual, defined using IC  [^3] Although the synthetic individual is defined by IC, the corresponding predictive favourable outcome probabilities calculation should follow Eq.
+(16)., can also be considered as a $\delta$-peer (The proof is given in Appendix C).
+Proposition 1.
+Let $A$ be an individual and $\mathcal{C}=\{C_{1},C_{2},\ldots\}$ denote all of $A$’s $\delta$-peers.
+Define a synthetic individual $\bar{T}_{i}$ using the average IC of any subset $\mathcal{C}_{i}$ of $K$ peers, where $\mathcal{C}_{i}=\{C_{1}^{i},\ldots,C_{K}^{i}\}\subseteq\mathcal{C}$, $i\in\{1,2,\ldots,N\}$ and $C_{j}^{i}$ represents the $j$-th peer in the $i$-th selection with the unprotected feature $\bm{x}_{j}^{i}$.
+This synthetic individual $\bar{T}_{i}=\sum_{j=1}^{K}\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}^{i})/K$ can also be considered as a $\delta$-peer of $A$.
+Consequently, by randomly selecting $K$ peers from the set of all observed $\delta$-peers $N$ times, we compute the predictive favourable outcome probabilities $\bar{T}_{i}$ for each $i$-th selection.
+We then use the mean of the resulting sample mean distribution, $\{\bar{T}_{i}\}_{i=1}^{N}$, consisting of all confirmed $\delta$-peers as per Proposition 1, to estimate the overall mean $\mu$ of favourable outcome probabilities across all peers.
+3.5 Peer-induced fairness auditing
+Finally, to formalise the process of auditing whether an individual in a protected group is subjected to algorithmic bias, we take advantage of hypothesis testing.
+This framework is predicated on an appropriate threshold for peer identification $\delta$ and a specific classifier $f$.
+It aims to test whether the sample mean distribution $\{\bar{T}_{i}\}_{i=1}^{N}$ is statistically equivalent to $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$.
+Since $\bar{T}_{i}$ follows a normal distribution and $N$ is a large enough number, our hypothesis is consistent with the standard $z$-test, which is designed to evaluate the presence of algorithmic bias statistically.
+  • $H_{0}$ (Null Hypothesis): The individual $A=(s_{-},\bm{x}_{0})$ is equally treated according to $(\delta,f)$-“peer-induced fairness” criterion,
+    $H_{0}:\mathbb{E}[\bar{T}_{i}]=\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0}).$ (17)
+  • $H_{1}$ (Alternative Hypothesis): The individual $A$ is subject to algorithmic bias under $(\delta,f)$-“peer-induced fairness” criterion, which is evidenced by a significant disparity in treatment compared to their unprotected peers,
+    $H_{1}:\mathbb{E}[\bar{T}_{i}]\neq\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0}).$ (18)
+Furthermore, it is also potential to consider two additional scenarios with one-sided tests: checking whether the individual is algorithmically discriminated against, where $H_{2}:\mathbb{E}[\bar{T}_{i}]<\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$, or algorithmically benefited, where $H_{3}:\mathbb{E}[\bar{T}_{i}]>\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$.
+3.6 Overall auditing workflow
+To illustrate the overall workflow of our “peer-induced fairness” tool, we present a flowchart in Figure 2 that visualises the steps involved in assessing potential algorithmic bias in an AI decision system.
+As an auditing tool, this framework can effectively determine whether the outcomes produced by an AI decision system exhibit algorithmic bias against a particular protected group.
+The process is straightforward and can function as a plug-and-play tool for not only AI developers but also regulators without access to the underlying decision process.
+Specifically, depending on the specific scenario, users can select different fitting and prediction models for computing $IC$ and predict $\mathbb{P}(\hat{Y}=1|s_{,}x)$ respectively for each instance in the datasets.
+Moreover, to handle varying characteristics of protected attributes and different levels of data quality, users have the flexibility to adjust the threshold $\delta$ in peer identification, which defines the degree of similarity required for an individual from a non-protected group to be considered a valid peer.
+This allows for a balance between the need for precise comparability and the practical constraints of the dataset.
+[…]
+Figure 2: The overall auditing workflow.
+“Compute $\mathbb{P}(\hat{Y}=1|s_{,}x)$” step requires a given prediction model, “Compute $IC$” step requires a given fitting model, “Find peers in B using $IC$, denote $\mathcal{C}(a)$” step requires a given $\delta$.
+Although the fitting model is usually the same as the prediction model, distinct choices are also allowed.
+A critical consideration in sampling is ensuring the use of the Central Limit Theorem, which typically requires at least 35 peers for each individual in the protected group and thus we can randomly select at least 30 from them.
+This should be taken into account when deciding on an appropriate threshold $\delta$.
+A threshold that is too high may result in too many distinct peers, while one that is too low may lead to too few peers, preventing further analysis.
+Additionally, the hypothesis testing of our framework is adaptable.
+Users can choose between one-sided or two-sided hypothesis tests, as discussed in Section 3.5, and set significance levels based on their specific needs.
+For a two-sided hypothesis test, the framework categorises results into “Fairly treated” and “Not fairly treated”.
+Advanced statistical tests can also be explored for more complex cases.
+While there might be some misunderstandings and ambiguities in audit results, particularly when individuals who are treated fairly by the system still receive unfavourable decisions.
+Our framework addresses these issues by providing explanations.
+Explanations are offered only for fairly treated individuals.
+This allows offer actionable advice on how they can achieve more favourable outcomes, significantly enhancing the customer service experience.
+Additionally, this analysis helps identify specific areas relevant institutions should pay attention to, such as particular features where these individuals may be under-performing.
+Conversely, for those who experience discrimination or privilege, the problem lies within the AI decision system, not with the applicants.
+In such cases, efforts should be concentrated on improving the AI system rather than providing suggestions to the applicants.
+Finally, as shown in Figure 2, the framework can also serve as a self-assessment tool before releasing an AI decision-making product.
+During internal testing, developers can easily integrate this framework to conduct various tests—such as using different prediction models, fitting models, and assessing different data qualities—even performing stress testing for algorithmic fairness.
+4 Experiment
+To demonstrate the effectiveness of our proposed “peer-induced fairness” framework in auditing algorithmic fairness, and to examine the current state of credit scoring concerning fairness, we experiment on the SMEs’ access to finance.
+The dataset used for this study is collected from the UK Archive Small and Medium-Sized Enterprise Finance Monitor (BDRC Continental,, 2023).
+The dataset compiles survey information on SMEs [^4] SMEs included in this survey meet the four criteria: 1) employ no more than 250 individuals, 2) have an annual turnover not exceeding £25 million, 3) do not operate as social enterprises or non-profit organisations, and 4) are not owned by another company by more than 50% (BDRC Continental,, 2023)., spanning from 2011Q1 to 2023Q4, with approximately 4,500 telephone interviews conducted per quarter across the UK.
+Each interview provides insights into the experiences of SMEs with external financing over the past 12 months, including their anticipated future financial needs and perceived obstacles to growth.
+It also details the characteristics of the SMEs and their owners or managers.
+Table 1: Description and abbreviation of features, grouped by whether they are intrinsic. The final column presents the values alongside their corresponding percentages.
+Features Abbreviation Value (Percentage)
+Non-intrinsic Features
+previous turn-down PT no (90.94%)
+yes (9.06%)
+finance qualification FQ no (45.66%)
+yes (54.34%)
+written plan WP no (37.58%)
+yes (62.42%)
+risk RI minimal (19.59%)
+low (43.11%)
+average (25.98%)
+above average (11.31%)
+product/service development PS no (70.25%)
+yes (29.75%)
+business innovation BI no (40.16%)
+yes (59.84%)
+loss or profit LP loss (86.07%)
+broken even (8.69%)
+profit (5.25%)
+turnover growth rate TG grown more than 20% (13.69%)
+grown but by less than 20% (40.33%)
+stayed the same (33.69%)
+declined (12.30%)
+funds injection FI no (67.17%)
+yes (32.83%)
+credit purchase CP no (18.48%)
+yes (81.52%)
+regular management account RM no (19.06%)
+yes (80.94%)
+Intrinsic Features
+principal PR construction (6.64%)
+agriculture, hunting and forestry (10.82%)
+fishing (12.01%)
+health and social work (12.62%)
+hotels and restaurants (11.68%)
+manufacturing (8.69%)
+real estate, renting and business activities (16.68%)
+transport, storage and communication (9.63%)
+wholesale/retail (11.23%)
+other community, social and personal service (9.63%)
+legal status LS sole proprietorship (4.88%)
+partnership (10.57%)
+limited liability partnership (7.50%)
+limited liability company (77.05%)
+startups SU no (97.5%)
+yes (2.5%)
+London & South East LS no (23.61%)
+yes (76.39%)
+We chose the SMEs dataset for two main reasons.
+First, SMEs play a crucial role in national economic development, making it vital for banks and financial institutions to provide essential support and ensure fair treatment.
+An algorithmic fairness auditing tool is, therefore, essential for financial regulators.
+It also serves as a self-assessment resource for lenders, helping them ensure compliance with the recent EU AI Act requirements in their AI decision systems during product development.
+However, due to limited access to SMEs’ data, there is a significant research gap regarding algorithmic bias in SMEs’ access to finance.
+Second, the SMEs dataset is survey-based and characterised by relatively low quality, making it an ideal choice for stress-testing our proposed framework to assess its effectiveness in handling such data challenges.
+To avoid redundancy, we selected survey results from 2012Q4 to 2020Q2 and focused on 15 important features identified in the literature (Sun et al.,, 2021; Calabrese et al.,, 2022; Cowling et al.,, 2016, 2022, 2012) (see Table 1 for details).
+These features capture various aspects of the loan application process.
+After filtering out data points with more than 20% missing features, the final dataset comprised 4,159 entries for analysis.
+Details of the data cleaning process can be found in Supplementary Materials.
+To apply our proposed framework, we consider the 15 features listed in Table 1 as $X$, and treat the firm size as the protected attribute $S$, defined by a combination of the number of employees and annual turnover (Micro-firms are defined as those with fewer than 10 employees and an annual turnover of less than £2 million following the literature (Sun et al.,, 2021)).
+Such grouping leads to 1,719 micro-firms ($s=s_{-}$) and 2,440 non-micro firms ($s=s_{+}$) as non-protected group.
+Our dataset does not exhibit a significant imbalance in the protected attributes, which is useful, as this would complicate testing our framework’s universality in different imbalance levels as in Section 5.2.
+Oversampling to adjust imbalance could alter feature relationships, making bias auditing unreliable (Chen et al.,, 2024).
+Instead, we maintain a moderate imbalance and vary the imbalance level by under-sampling.
+For the target variable, we use the outcome of bank loan application, due to the significant role of bank loans in SME financing (Sun et al.,, 2021).
+The dataset records 3,391 approvals ($y=1$) and 768 rejections ($y=0$), highlighting the decisions faced by SMEs in access to finance.
+Following the workflow outlined in Figure 2, we focus on the 1,719 micro-firms to determine whether they have experienced algorithmic bias.
+As described in the workflow, we use logistic regression as the default model for both prediction and fitting for simplicity (see performance evaluation and robustness tests in Supplementary Materials).
+For the fitting and prediction, the data are typically split into training (80%) and testing (20%) sets, with hyper-parameters optimised via grid search and 5-fold cross-validation.
+The model yielding the highest AUC value is selected for predictions on the target $Y$.
+Without loss of generality, we set the default $\delta$ to 0.3 times the standard deviation of the micro-firms’ ICs.
+This flexible threshold can be adjusted according to the specific dataset and research context.
+Additional robustness tests regarding threshold adjustments are provided in Supplementary Materials.
+The remaining settings are $N=100$ and $K=30$.
+We include only micro-firms with more than 35 peers to meet the basic requirement for a large sample size.
+Due to the limitations of our dataset quality, firms with fewer than 35 peers are labelled as “Unknown” and, for illustrative purposes, will not be included in our further analysis.
+However, for real auditing tasks, the dataset size and quality are typically higher than those obtained from surveys, and this issue is likely to be mitigated.
+For the hypothesis testing step, the default test is $H_{0}$ vs.
+$H_{1}$ with a significance level of 5%.
+However, to differentiate between cases of discrimination and privilege, we also run tests for $H_{2}$ and $H_{3}$ against their respective alternative hypotheses.
+These tests compare the mean approval likelihood of the peers against that of the micro-firms, thereby identifying potential algorithmic bias in terms of discrimination or privilege.
+5 Results
+In this section, we present the experimental results on the SMEs dataset, demonstrating the efficacy of our “peer-induced fairness” framework.
+5.1 Algorithmic fairness auditing
+Following the workflow outlined in Figure 2 and the experimental settings described in Section 4, we successfully identified algorithmic bias within the SMEs dataset.
+The scatter plot in Figure 3, which compares approval likelihoods between micro-firms and their peers, reveals that only 2.48% of micro-firms are treated fairly, indicating significant disparities in the credit approval system.
+The remaining 97.52% experience algorithmic bias, with 41.51% of micro-firms facing discrimination.
+Interestingly, 56.40% of micro-firms, despite being underrepresented, benefit from the decision system by receiving approval likelihoods higher than the average of their peers.
+Figure 3: Comparative analysis of loan approval likelihood for micro-firms against peers.
+The black dashed 45-degree line, denoting $Y=X$, symbolises perfect fairness.
+Red and orange data points represent micro-firms with approval likelihoods significantly lower or higher, respectively than the average of their peers.
+Blue points denote no significant difference.
+Figure 4: Comparative analysis of loan approval likelihood for micro-firms under each algorithmic treatment category against peers.
+(a)-(c): Extremely discriminated (ED) micro-firms.
+(d)-(f): Fairly treated (FT) micro-firms.
+(g)-(i): Extremely privileged (EP) micro-firms.
+The coloured data points in the first column of each row represent a comparison among peers within each category.
+The x-axis shows the approval likelihood for micro-firms, while the y-axis displays the average approval likelihood of the peers.
+The second column compares the approval likelihood between these micro-firms (i.e., coloured points in (a), (d), (g)) and their peers at the group level.
+The third column provides the comparison, at the individual level, between the selected micro-firm (i.e., coloured triangle in (a), (d), (g)) and its peers.
+To identify the specific extent of discrimination and privilege faced by each micro-firm, we compare the approval likelihood difference between a given micro-firm $A=(s-,\bm{x_{0}})$ and its peers.
+For micro-firms with a higher likelihood of approval, we allow for greater tolerance when assessing extreme algorithmic bias, adjusting the standard based on each firm’s approval likelihood.
+Specifically, we consider a micro-firm to experience extreme algorithmic bias if the absolute difference exceeds 0.1 times its own approval likelihood.
+Mathematically, this is expressed as $|\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})-\mathbb{E}[\bar{T}_{i}]|>0.1\times\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$.
+A negative difference indicates discrimination, while a positive difference signifies privilege.
+This approach ensures the flexibility of the standard, making it suitable for firms in different situations.
+Instead, if the absolute difference is less than the threshold, it represents slight discrimination or slight privilege.
+Further details can be found in the Supplementary Materials.
+In certain cases, such slightly unfair treatment may be considered fair, depending on the regulatory tolerance and the specific industry being audited.
+Specifically, in our case, 26.71% of micro-firms experience extreme discrimination, with their approval likelihood markedly lower than that of their peers, as shown in Figure 4 (a)-(c), at both group and individual levels.
+32.17% of micro-firms are extremely privileged, as shown in Figure 4 (g)-(i).
+Even though algorithmic privilege might seem beneficial for micro-firms, neither scenario is desirable.
+We advocate for transparency and fairness in decision-making processes.
+Arbitrary or opaque factors influencing decisions are contrary to the principles of fairness and should be rigorously addressed to ensure equitable treatment across all applicants.
+It is important to emphasise that our framework is a tool for audits by regulators and stakeholders, aiming to detect algorithmic bias.
+In credit loan applications, rejected customers are particularly concerned about whether they were rejected and discriminated against, while regulators and banks require detailed results to audit the fairness of their models for all applicants.
+Therefore, our framework also includes detailed information on accepted applicants.
+Additionally, without compromising generalisation to other research areas, it is crucial to focus on all applicants.
+We also validate our framework by investigating the connection between accessing finance outcomes and disparities in algorithmic bias.
+Among these markedly discriminated micro-firms, 52.42% were denied loans, whereas only 9.97% of their peers faced rejection, highlighting a significant disparity in rejection rates.
+The rejection rate of micro firms decreases and that of their peers increases with the diminished discrimination.
+The difference in rejection rates between micro-firms and their peers also decreases.
+The rejection rates of peers fluctuate around the rejection rate of fairly treated micro-firms.
+This fluctuation indicates that within the category, some micro-firms experience higher rejection rates compared to their peers, while others experience lower rejection rates, illustrating a gradual convergence in rejection rates across categories with less pronounced discrimination.
+Notably, even the lowest peer rejection rate at the bottom of the error bar surpasses that of micro-firms in the extremely privileged category, where micro-firms experience the lowest rejection rates, as in Figure 5.
+These findings, derived from our bias audit based on financing outcomes prediction, align with the observed financing results.
+This congruence further validates the utility of our framework in accurately reflecting disparities and biases in the loan approval process.
+Further details on the extent of algorithmic bias are provided in Supplementary Materials, which expands the analysis to include two additional categories: slightly discriminated and slightly privileged.
+The analysis shows that even with these detailed treatment categories, the results consistently validate the effectiveness of our framework.
+Figure 5: Rejection rates of micro-firms across algorithmic treatment categories and their peers.
+The algorithmic treatment categories include extremely discriminated (ED), fairly treated (FT), and extremely privileged (EP).
+Each category includes multiple micro-firms with a single rejection rate, shown as histograms, while the rejection rate of peers of each micro-firm in this category is represented in the black line with error bars to indicate variability.
+The above experimental results reveal that our “peer-induced fairness” framework not only effectively identifies disparities in algorithmic fairness but also facilitates the visual representation of individual-level discrepancies across all users in the dataset.
+This capability enables clear visualisation of algorithmic fairness, making discrimination or privilege readily distinguishable.
+Such insights are invaluable for both regulatory purposes and for verifying the effectiveness of algorithmic fairness models.
+5.2 Data scarcity and imbalance
+Data scarcity and imbalance significantly influence the performance of advanced machine learning models due to the potential for inaccurate parameter estimation (Wang et al.,, 2024).
+This issue is especially pronounced in many datasets, where the representation of minority groups is often limited compared to majority groups (Chen et al.,, 2024; Lessmann et al.,, 2015).
+This discrepancy caused by the poor data quality, subsequently affects the auditing of algorithmic bias.
+Our “peer-induced fairness” framework addresses these challenges uniquely.
+Unlike traditional models that rely heavily on the data from the protected group, our framework bases all parameter estimations on peers identified within the unprotected group.
+This group typically possesses ample data points, effectively mitigating issues related to data scarcity and group imbalance, making our framework robust theoretically.
+We investigate the stability and credibility of our “peer-induced fairness” framework by evaluating the percentages of unfairly treated ($PUT$) protected individuals or organisations and the invariant outcome ratio ($IOR$) under varying levels of imbalance.
+The imbalance ratio, $\omega$, is defined as the proportion of samples in the protected class:
+$\omega=\frac{\#(S=s_{-})}{\#(S=s_{+})+\#(S=s_{-})},$ (19)
+where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality of a set.
+A perfectly balanced dataset corresponds to $\omega=50\%$.
+The $PUT$ is calculated as the number of unfairly treated individuals or organisations divided by the total number of selected subjects in the experiments with different $\omega$.
+The $IOR$ is computed as the number of selected individuals or organisations in the experiment with $\omega$ that have unchanged predictive outcomes compared to the original experiment divided by the number of commonly selected subjects in both the experiment with $\omega$ and the original experiment.
+In the SMEs experiment, building upon the default settings outlined in Section 4, we investigate the impact of varying imbalance ratios by randomly selecting subsets of the original dataset with controlled imbalance levels.
+Specifically, we evaluate the framework’s performance at imbalance ratios of $\omega=\{36.33\%,31.33\%,26.33\%,21.33\%,16.33\%,11.33\%\}$, where the original dataset’s imbalance ratio is $\omega=41.33\%$.
+By gradually decreasing the proportion of micro-firms in these subsets, we assess the framework’s robustness across different levels of imbalance.
+To minimise the effects of randomness in subset selection, this process is repeated five times.
+The detailed procedure is provided in Supplementary Materials.
+The results are visualised in Figure 6 and demonstrate the robustness of our framework.
+From the view of $PUT$, the small error bars across all the imbalance levels suggest the results across the five repetitions are highly consistent.
+This observation underscores the robustness of our “peer-induced framework” to imbalanced datasets.
+From the view of $IOR$, it is approximately 95% and remains stable across different imbalance levels.
+This aligns with our expectations, as the framework does not rely on data from the minority group but rather leverages information from the unprotected group, leading to inherent robustness.
+The small error bars also suggest that the results regarding $IOR$ in these five repeats are highly consistent.
+These findings highlight the universality of our “peer-induced fairness” framework with respect to the different data quality, as the auditing results remain consistent despite variations in imbalance levels.
+This distinguishes our framework from others by effectively addressing the prevalent challenges of data scarcity and imbalance in the field.
+Regulators can utilise this framework to evaluate the practices of companies and institutions, while these organisations can also reliably employ it for thorough self-assessment.
+Additionally, an alternative computation method is detailed in Supplementary Materials to further enhance the robustness of our approach.
+Figure 6: Percentage of unfairly treated micro-firms and invariant outcome ratio at different group imbalance levels.
+The imbalance level is represented on the x-axis as a percentage, ranging from 11.33% to 36.33%.
+The left y-axis shows the percentage of unfairly treated micro-firms (blue line), while the right y-axis displays the invariant outcome ratio (red line) as the imbalance level changes from the initial level to other levels.
+5.3 Explainable fairness discovery
+Next, we focus on the final step outlined in Figure 2, which involves providing explanations for individuals who are fairly treated but receive an outcome of $y=0$.
+Our explanation approach is based on comparing the features of these individuals with those of their peers.
+This method helps avoid misunderstandings and ambiguity, allowing us to offer a clear “watch-out” list of features that may have contributed to the unfavourable decision.
+Figure 7: Comparative analysis of non-intrinsic features for rejected while fairly treated micro-firms vs.
+accepted peers.
+The x-axis represents the selected key attributes being analysed, including finance qualification for manager (FQ), written plan (WP), previous turn-down (PT), loss or profit (LP), risk (RI), product/service development (PS), business innovation (BI), regular management account (MA), turnover growth rate (TG), credit purchase (CP) and funds injection (FI).
+The y-axis represents the percentage of those micro-firms with significantly worse performance than their accepted peers on each feature.
+Given the existence of accepted peers as the counterfactual instances with positive accessing finance outcomes, the micro-firm which is fairly treated should originally have the same outcomes.
+Our framework identifies the feature differences between each rejected while fairly treated micro-firm and its accepted peers by another hypothesis testing, with details presented in Supplementary Materials.
+For each feature, we summarise the percentage of these micro-firms that perform significantly worse than their accepted peers.
+We consider some non-intrinsic features to identify and understand these discrepancies, as in Figure 7.
+The descriptions for each feature value are shown in Supplementary Materials.
+Results show that even though none of them have been rejected previously and only 25% of them perform worse on financial qualifications and written plans, banks generally prioritise the financial and business health of firms.
+75% of these micro-firms invest excessively in business innovation and have lower risk ratings.
+Besides, half of them invest in product/service development and have lower profits.
+The uncertain returns and high risks associated with innovation lead to the failure or commercial non-viability of most innovative products (Coad and Rao,, 2008; Hall,, 2002; Freel,, 2007), exacerbating already poor-performing risk indicators.
+The worse performance on these key features makes banks cautious about the long-term financial sustainability of these firms.
+It also reflects the capability of these micro-firms, negatively affecting their loan approvals.
+This exploration identifies the differences between micro-firms and their peers for each feature and summarises the percentage of micro-firms that perform worse on each feature.
+This explainable analysis not only enhances the transparency of our framework but also supports regulators and stakeholders in understanding the specific challenges most incapable micro-firms face, and highlights the features that they need to watch out for and pay extra attention to.
+6 Concluding remarks and discussion
+In the age of AI, where automated decision-making systems increasingly determine access to essential services such as finance, housing, and employment, the consequences of algorithmic bias can be severe.
+Discriminatory practices not only undermine social equity but also violate legal and ethical standards, potentially causing significant harm to vulnerable groups.
+Various regulatory documents have emphasised the need for ongoing oversight and auditing of decision systems (Voigt and Von Dem Bussche,, 2017; Madiega,, 2024; British Standards Institution,, 2023), both at the initial deployment stage and throughout their operational life cycle (Madiega,, 2024).
+This focus on fairness is not confined to the European Union; several countries, including the United Kingdom and the United States (President and Press,, 2016; British Standards Institution,, 2023), have also introduced regulations and guidelines to ensure that AI and automated decision-making systems function transparently and equitably.
+To meet the growing regulatory demands across, we proactively and timely introduce an algorithmic fairness auditing framework.
+It is a robust auditing framework for both internal and external assessment in a plug-and-play fashion.
+Designed as a fully modular tool, the framework allows users—whether financial institutions, regulators, or third-party auditors—to customise settings based on their specific needs and objectives, making it highly adaptable across various sectors.
+The core idea is grounded in peer comparisons, which is both intuitive and intrinsic.
+This approach is computationally efficient and robust to varying data qualities, ensuring reliable auditing results in different scenarios.
+By facilitating comprehensive fairness audits, our tool helps prevent automated systems from perpetuating or exacerbating existing inequalities.
+Our framework also enhances transparency by providing necessary explanations and “watch-out” lists for those who receive unfavourable decisions due to insufficient capabilities.
+This feature promotes understanding and trust among users and affected groups, aligning with regulatory requirements for transparency and accountability.
+In a regulatory landscape where continuous monitoring and auditing are increasingly mandated, such tools are indispensable.
+They offer a practical means to ensure that AI systems are both legally compliant and socially responsible, adhering to broader ethical imperatives for fairness and accountability.
+From an empirical standpoint, our framework uncovers alarming issues in the current state of SMEs’ access to finance.
+Specifically, our findings indicate that only 2.48% of micro-firms are treated fairly, while a staggering 41.51% face discrimination.
+Even when we adjusted the data quality by altering the imbalance level, the audit results remained highly consistent with our original findings.
+These results underscore a serious and inequitable banking environment that demands immediate attention.
+Additionally, we observed that some micro-firms are rejected due to inherent limitations, such as higher risk or greater investment in innovation, rather than discrimination when compared to their peers.
+These empirical findings also demonstrate the effectiveness and robustness of our framework in real-world applications.
+For these reasons, we believe our framework represents a significant advancement and a policy-relevant contribution to algorithmic fairness.
+Given the modular structure of our framework, there is significant potential for further enhancement and adaptation.
+Currently, our framework is based on a static causal model, which, while effective for many applications, may not fully capture the complexities of real-world scenarios where dynamic causal models are more appropriate.
+In such cases, feedback loops can alter relationships over time, as decisions made based on certain features can influence future data and outcomes.
+A static framework may not adequately account for these evolving interactions.
+However, the core concept of peer comparison remains valid even within a dynamic causal model.
+Future studies could focus on integrating dynamic causal modelling into our framework to better address these feedback mechanisms, ensuring its applicability and robustness across a broader range of contexts.
+Data availability
+Data and codes are available at UK Data Archive and GitHub respectively.
+Acknowledgements
+The authors of this manuscript would like to thank Prof.Raffaella Calabrese and Dr.Yizhe Dong for their assistance and support in the discussion and research direction.
+Appendix Appendix A Proof of Theorem 1
 Proof.
@@
 ∎
-Consequently, we can generate an algorithm shown in Algorithm 1 to identify all peers in the dataset.
-Input: A set of individuals $\{A\}=\{(s_{-},\bm{x}_{0})\}$ from the protected group, a set of individuals $\{B_{i}\}_{i=1}^{N}=\{(s_{+},\bm{x}_{i})\}$ from the unprotected group, a threshold $\delta$, and a minimum number of peers $U$.
-Output: A subset of $\{B_{i}\}$ designated as $\delta$-peers of $A$, with each protected individual having at least $U$ peers.
-1 foreach $A=(s_{-},\bm{x}_{0})$ do
-2       Initialise an empty list of peers for $A$, denoted as $\text{Peers}_{A}$
-3       Compute $\xi(s_{-},\bm{x}_{0})$ for $A$
-4       foreach $B_{i}=(s_{+},\bm{x}_{i})$ in $\{B_{i}\}_{i=1}^{N}$ do
-5             Compute $\xi(s_{+},\bm{x}_{i})$ for $B_{i}$
-6             Calculate the difference $\Delta=|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{i})|$
-7             if $\Delta<\delta$ then
-8                   Add $B_{i}$ to $\text{Peers}_{A}$
-9
-10             end if
-11
-12       end foreach
-13
-14 end foreach
-Algorithm 1 Identification of $\delta$-Peers for Protected Group Individuals
-4.4 Peer-induced fairness
-Following the idea of peer comparison, definition, and identification, we can now introduce the concept of peer-induced fairness.
-Definition 4 ($(\delta,f)$-peer-induced fairness [^2] Although the term “peer-induced fairness” has been used in other contexts as noted by (Ho and Su,, 2009; Li and Jain,, 2016), our concept is novel in its reliance on a structured causal reasoning framework specifically tailored for classification tasks.).
-Consider an individual $A=(s_{-},\bm{x}_{0})$ and assuming $A$ has a number of $\delta$-peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ where $C_{j}=(s_{+},\bm{x}_{j})$.
-$A$ is said to be fairly treated by the peers subject to $(\delta,f)$ if and only if
-$\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})=\mathbb{E}_{C_{j}\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|C_{j})],$ (16)
-where $\hat{Y}$ is the predictive outcome provided with the classifier $f$.
-As discussed in previous sections, while we can directly estimate $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$ from individual observations, estimating the expected value $\mathbb{E}_{C_{j}\in\mathcal{C}}[\mathbb{P}(\hat{Y}_{s_{+}}|C_{j})]$ presents challenges due to the limited number of observations available for $\delta$-peers.
-Consequently, we have to rely on observable peers to approximate the population mean.
-To formalise this, we introduce the random variable
-$T_{j}=\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}).$ (17)
-Upon examining the distribution of $T_{j}$, we find that it does not follow a normal distribution, with details presented in Supplementary Materials.
-Therefore, we randomly select a subset of peers and use the sample mean to estimate the population mean,
-$\bar{T}=\frac{1}{K}\sum_{j=1}^{K}\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}),$ (18)
-where $K$ is a large enough number of peers in the subset.
-According to the Central Limit Theorem, the sample mean $\bar{T}$ follows a normal distribution, and thus $\mathbb{E}[\bar{T}]$, can be employed to estimate the overall predictive probabilities of favourable outcomes among peers, denoted as $\mathbb{E}[T]=\mu$.
-Based on this, we propose a proposition that a synthetic individual, defined using $IC$  [^3] Although the synthetic individual is defined by $IC$, the corresponding predictive favourable outcome probabilities calculation should follow Eq.
-(18)., can also be considered as a $\delta$-peer.
-Proposition 1.
-Let $A$ be an individual and $\mathcal{C}=\{C_{1},C_{2},\ldots\}$ denote all of $A$’s $\delta$-peers.
-Define a synthetic individual $\bar{T}_{i}$ using the average $IC$ of any subset $\mathcal{C}_{i}$ of $K$ peers, where $\mathcal{C}_{i}=\{C_{1}^{i},\ldots,C_{K}^{i}\}\subseteq\mathcal{C}$, $i\in\{1,2,\ldots,N\}$ and $C_{j}^{i}$ represents the $j$-th peer in the $i$-th selection with the unprotected feature $\bm{x}_{j}^{i}$.
-This synthetic individual $\bar{T}_{i}$ can also be considered as a $\delta$-peer of $A$.
+Appendix Appendix B Implementation for peer identification
+Algorithm 1 Identification of $\delta$-Peers for Protected Individuals
+1:A set of individuals $\{A\}=\{(s_{-},\bm{x}_{0})\}$ from the protected group, a set of individuals $\{B_{i}\}_{i=1}^{N}=\{(s_{+},\bm{x}_{i})\}$ from the unprotected group, a threshold $\delta$, and a minimum number of peers $U$.
+2:A subset of $\{B_{i}\}$ designated as $\delta$-peers of $A$, with each protected individual having at least $U$ peers.
+3:for all $A=(s_{-},\bm{x}_{0})$ do
+4:     Initialise an empty list of peers for $A$, denoted as $\text{Peers}_{A}$
+5:     Compute $\xi(s_{-},\bm{x}_{0})$ for $A$
+6:     for all $B_{i}=(s_{+},\bm{x}_{i})$ in $\{B_{i}\}_{i=1}^{N}$ do
+7:         Compute $\xi(s_{+},\bm{x}_{i})$ for $B_{i}$
+8:         Calculate the difference $\Delta=|\xi(s_{-},\bm{x}_{0})-\xi(s_{+},\bm{x}_{i})|$
+9:         if $\Delta<\delta$ then
+10:              Add $B_{i}$ to $\text{Peers}_{A}$
+11:         end if
+12:     end for
+13:end for
+Appendix Appendix C Proof of Proposition 1
 Proof.
@@
 ∎
-Consequently, by randomly selecting $K$ peers from the set of all observed $\delta$-peers $N$ times, we calculate the predictive favourable outcome probabilities $\bar{T}_{i}$ for each $i$-th selection using Eq.
-(18).
-Here
-$\bar{T}_{i}=\frac{1}{K}\sum_{j=1}^{K}\mathbb{P}(\hat{Y}_{s_{+}}|C_{j}^{i}).$ (19)
-We then utilise the mean of the observed sample mean distribution, $\{\bar{T}_{i}\}_{i=1}^{N}$, which are all confirmed $\delta$-peers as per Proposition 1, to estimate the overall mean $\mu$ of favourable outcome probabilities among all peers.
-4.5 Hypothesis testing for peer-induced fairness
-Finally, to formalise the process of auditing whether an individual in a protected group is subjected to algorithmic bias, we propose a hypothesis-testing framework.
-This framework is predicated on an appropriate threshold for peer identification $\delta$ and a specific classifier $f$.
-It aims to test whether the sample mean distribution $\{\bar{T}_{i}\}_{i=1}^{N}$ is statistically equivalent to $\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x})$.
-Since $\bar{T}_{i}$ follows a normal distribution and $N$ is a large enough number, our hypothesis is consistent with the standard $z$-test, which is designed to evaluate the presence of algorithmic bias statistically.
-  • $H_{0}$ (Null Hypothesis): The individual $A=(s_{-},\bm{x}_{0})$ is equally treated according to $(\delta,f)$-“peer-induced fairness” criterion,
-    $H_{0}:\mathbb{E}[\bar{T}_{i}]=\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0}).$ (20)
-  • $H_{1}$ (Alternative Hypothesis): The individual $A$ is subject to algorithmic bias under $(\delta,f)$-“peer-induced fairness” criterion, which is evidenced by a significant disparity in treatment compared to their unprotected peers,
-    $H_{1}:\mathbb{E}[\bar{T}_{i}]\neq\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0}).$ (21)
-Furthermore, it is also feasible to consider two additional scenarios: checking whether the individual is algorithmically discriminated against, where $H_{2}:\mathbb{E}[\bar{T}_{i}]<\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$, or algorithmically benefited, where $H_{3}:\mathbb{E}[\bar{T}_{i}]>\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$.
-5 Experiment setup
-Fairness concepts apply not only to decision-making concerning groups of people but also to broader domains, such as companies within the economic system.
-The banking loan approval algorithm that uses historical data often places micro-firms at a disadvantage due to their smaller size and limited historical records compared to larger firms (Cenni et al.,, 2015).
-Consequently, these smaller entities may face higher interest rates or be more likely to be denied loans, despite their growth potential.
-Therefore, to illustrate the usefulness of our novel “peer-induced fairness” framework, we consider a real example using SMEs data.
-The collected data is from the UK Archive Small and Medium-Sized Enterprise Finance Monitor (BDRC Continental,, 2023).
-The dataset compiles survey information on SMEs  [^4] SMEs included in this survey meet the four criteria: 1) employ no more than 250 individuals, 2) have an annual turnover not exceeding £25 million, 3) do not operate as social enterprises or non-profit organisations, and 4) are not owned by another company by more than 50%.
-spanning from 2011Q1 to 2023Q4, with approximately 4,500 telephone interviews conducted per quarter across the UK.
-Each interview provides insights into the experiences of SMEs with external financing over the past 12 months, including their anticipated future financial needs and perceived obstacles to growth.
-It also details the characteristics of the SMEs and their owners or managers.
-To avoid redundancy, we selected survey results from 2012Q4 to 2020Q2.
-We focused on 15 important features identified from the literature (Sun et al.,, 2021; Calabrese et al.,, 2022; Cowling et al.,, 2016, 2022, 2012) (refer to Table 1 for details).
-These features demonstrate various dimensions of the loan application process.
-After filtering out data points with more than 20% missing features, we obtained a dataset comprising 4,159 data points for our analysis.
-The details of data cleaning are presented in Supplementary Materials.
-Table 1: Description of features.
-Feature Category Value Percentage
-risk multivariate & ordinal minimal 19.59%
-low 43.11%
-average 25.98%
-above average 11.31%
-principal multivariate & nominal construction 6.64%
-agriculture, hunting and forestry 10.82%
-fishing 12.01%
-health and social work 12.62%
-hotels and restaurants 11.68%
-manufacturing 8.69%
-real estate, renting and business activities 16.68%
-transport, storage and communication 9.63%
-wholesale/retail 11.23%
-other community, social and personal service 9.63%
-legal status multivariate & nominal sole proprietorship 4.88%
-partnership 10.57%
-limited liability partnership 7.50%
-limited liability company 77.05%
-loss or profit multivariate & ordinal loss 86.07%
-broken even 8.69%
-profit 5.25%
-turnover growth rate multivariate & ordinal grown more than 20% 13.69%
-grown but by less than 20% 40.33%
-stayed the same 33.69%
-declined 12.30%
-funds injection binary no 67.17%
-yes 32.83%
-credit purchase binary no 18.48%
-yes 81.52%
-startups binary no 97.5%
-yes 2.5%
-previous turn-down binary no 90.94%
-yes 9.06%
-London & South East binary yes 76.39%
-no 23.61%
-business innovation binary no 40.16%
-yes 59.84%
-product/service development binary no 70.25%
-yes 29.75%
-regular management account binary no 19.06%
-yes 80.94%
-written plan binary no 37.58%
-yes 62.42%
-finance qualification binary no 45.66%
-yes 54.34%
-For this analysis, we designate the observed features listed in Table 1 as $X$.
-We treat the firm size as the protected attribute $S$, defined by a combination of the number of employees and annual turnover.
-This categorisation results in 1,719 micro-firms ($s=s_{-}$) classified as protected and 2,440 non-micro firms ($s=s_{+}$) as unprotected [^5] Micro-firms are defined as those with fewer than 10 employees and an annual turnover of less than £2 million (Sun et al.,, 2021)..
-We consider the outcome of bank loan applications, denoted as $Y$, due to the significant role of bank loans in SME financing (Sun et al.,, 2021).
-The dataset records 3,391 approvals ($y=1$) and 768 rejections ($y=0$), highlighting the decisions faced by SMEs in securing financial support.
-In subsequent analyses, we focus on the 1,719 micro-firms to determine if they have experienced algorithmic bias using our “peer-induced fairness” framework.
-Initially, it is essential to identify each firm’s peers by computing the $IC$ as specified in Algorithm 1.
-Without loss of generality, we set the default $\delta$ to 0.3 times the standard deviation of the micro-firms $ICs$.
-This flexible threshold can be adjusted based on the dataset of the specific research field.
-The specific robustness tests are presented in Supplementary Materials.
-However, direct estimation of $IC$ from observed data is challenging, necessitating the use of a fitted model.
-For simplicity, we employ a logistic classifier to estimate the probability of an individual being labelled as part of the protected group, with performance and robustness test in Supplementary Materials.
-Following Proposition 1, we then utilise the expectation of the observed sample mean distribution to estimate the overall mean of all peers.
-As a standard approach, we randomly sample $N=100$ times, and each sample consists of $K=30$ data points.
-We consider only those micro-firms with more than $U=35$ peers; firms with fewer than 35 peers are labelled as “Unknown”.
-Again, direct estimation of the mean from Eq.
-(18) is not feasible, requiring the use of a predictive classification model.
-By default, we use a logistic classifier with performance in Supplementary Materials, although any classification model could be substituted.
-The robustness of the classification model is demonstrated in Supplementary Materials.
-The data are typically split into training (80%) and testing (20%) sets, with hyper-parameters optimised via grid search and 5-fold cross-validation.
-The model yielding the highest AUC value is selected for predictions on the target $Y$.
-Ultimately, we conduct hypothesis tests (i.e., $H_{0}$, $H_{2}$, $H_{3}$) to compare the mean approval likelihood of the peers against that of the micro-firms, thereby identifying potential algorithmic bias, discrimination and privilege.
-The significance level for the hypothesis test is set at $\alpha=5\%$.
-Due to the presence of “Unknown” micro-firms, hypothesis testing is performed on a subset of the data, consisting of 1007 data points.
-6 Experiment results
-In this section, we present the experimental results derived from the SMEs data to demonstrate the efficacy of our “peer-induced fairness” framework.
-These results illustrate how the framework operates in practice and underscore its potential for auditing bias in algorithmic decision-making.
-6.1 Algorithmic fairness auditing
-With the evolution of algorithmic fairness methods and the increasing regulatory demands for data protection and transparency in decision-making processes, there is a growing emphasis on applying advanced methodologies to ensure fairness in decision systems.
-For instance, in the banking sector, it is becoming common to self-scrutinise or externally audit decision-making processes to determine whether they meet established fairness criteria or if they continue to exhibit significant algorithmic bias.
-Following the steps outlined in Section 5, we have identified algorithmic bias within this SMEs dataset.
-The scatter plot (refer to Fig.
-2) comparing micro-firms to their peers regarding approval likelihood reveals that only 2.48% of them are treated fairly, indicating significant disparities in the credit approval system.
-97.52% of them experience algorithmic bias, with 41.51% of micro-firms experiencing discrimination.
-An intriguing observation is that the remaining 56.40% of micro-firms, despite being under-represented, benefit from the decision system, receiving approval likelihoods higher than the average for their peers.
-Figure 2: Comparative analysis of loan approval likelihood for micro-firms against peers.
-The black dashed 45-degree line, denoting $Y=X$, symbolises perfect fairness.
-Red and orange data points represent micro-firms with approval likelihoods significantly lower or higher, respectively than the average of their peers.
-Blue points denote no significant difference.
-To identify the specific extent of discrimination and privilege faced by each micro-firm, we compare the approval likelihood difference between a given micro-firm $A=(s-,\bm{x_{0}})$ and its peers.
-For micro-firms with a higher likelihood of approval, we allow for greater tolerance when assessing extreme algorithmic bias, adjusting the standard based on each firm’s approval likelihood.
-Specifically, we consider a micro-firm to experience extreme algorithmic bias if the absolute difference exceeds 0.1 times its own approval likelihood.
-Mathematically, this is expressed as $|\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})-\mathbb{E}[\bar{T}_{i}]|>0.1\times\mathbb{P}(\hat{Y}_{s_{-}}|s_{-},\bm{x}_{0})$.
-A negative difference indicates discrimination, while a positive difference signifies privilege.
-This approach ensures the flexibility of the standard, making it suitable for firms in different situations.
-Specifically, 26.71% of micro-firms experience substantial discrimination, with their approval likelihood markedly lower than that of their peers, as shown in Panel (a)-(c) in Fig.
-3, at both group and individual levels.
-32.17% of micro-firms are extremely privileged, as shown in Panel (g)-(i).
-Even though algorithmic privilege might seem beneficial for micro-firms, neither scenario is desirable.
-We advocate for transparency and fairness in decision-making processes.
-Arbitrary or opaque factors influencing decisions are contrary to the principles of fairness and should be rigorously addressed to ensure equitable treatment across all applicants.
-Figure 3: Comparative analysis of loan approval likelihood for micro-firms under each algorithmic treatment category against peers.
-(a)-(c): Extremely discriminated (ED) micro-firms.
-(d)-(f): Fairly treated (FT) micro-firms.
-(g)-(i): Extremely privileged (EP) micro-firms.
-The second column provides the approval likelihood comparison between these micro-firms and their peers at the group level.
-The third column provides the comparison, at the individual level, between the selected micro-firm in the first column and its peers.
-It is important to emphasise that our framework is a tool for audits by regulators and stakeholders, aiming to detect algorithmic bias.
-In the credit loan application, rejected customers are particularly concerned about whether they were treated fairly, while regulators and banks require detailed results to audit the fairness of their models.
-Therefore, our framework also includes detailed information on accepted applicants.
-Additionally, without compromising generalisability to other research areas, it is crucial to focus on all applicants.
-We also validate our framework by investigating the connection between accessing finance outcomes and disparities in algorithmic bias.
-Among these markedly discriminated micro-firms, 52.42% were denied loans, whereas only 9.97% of their peers faced rejection, highlighting a significant disparity in rejection rates.
-The rejection rate of micro firms decreases and that of their peers increases with the diminished discrimination.
-The difference in rejection rates between micro-firms and their peers also decreases.
-The rejection rates of peers fluctuate around the rejection rate of fairly treated micro-firms.
-This fluctuation indicates that within the category, some micro-firms experience higher rejection rates compared to their peers, while others experience lower rejection rates, illustrating a gradual convergence in rejection rates across categories with less pronounced discrimination.
-Notably, even the lowest peer rejection rate surpasses that of micro-firms in the extremely privileged category, where micro-firms experience the lowest rejection rates, as in Fig.
-4.
-These findings, derived from our bias audit based on financing outcomes prediction, align with the observed financing results.
-This congruence further validates the utility of our framework in accurately reflecting disparities and biases in the loan approval process.
-Further details of the degree of algorithmic bias are provided in Supplementary Materials.
-Figure 4: Rejection rates of micro-firms across algorithmic treatment categories and their peers.
-The algorithmic treatment categories include extremely discriminated (ED), fairly treated (FT), and extremely privileged (EP).
-Each category includes multiple micro-firms with a single rejection rate, shown as histograms, while the rejection rate of peers of each micro-firm in this category is represented in the black line with error bars to indicate variability.
-From the analysis presented, it is evident that our “peer-induced fairness” framework not only identifies disparities in algorithmic fairness but also facilitates the visual representation of individual-level discrepancies across all users in the dataset.
-This capability allows for clear visualisation of algorithmic fairness, where discrimination or benefit is readily distinguishable.
-Such insights are invaluable not only for regulatory purposes but also for verifying the effectiveness of algorithmic fairness models.
-Furthermore, we subjected all results to robustness tests, varying the level of peer identification threshold, model fitting selection, and prediction algorithms to ensure the integrity of our findings.
-6.2 Data scarcity and imbalance
-Data scarcity and imbalance significantly influence the performance of advanced machine learning models due to the potential for inaccurate parameter estimation (Chen et al.,, 2024; Lessmann et al.,, 2015).
-This issue is particularly pronounced in the field of algorithmic fairness, where the representation of minority groups is often limited compared to majority groups.
-This discrepancy caused by the poor data quality, subsequently affects the assessment of algorithmic fairness.
-Our “peer-induced fairness” framework addresses these challenges uniquely.
-Unlike traditional models that rely heavily on the data from the protected group, our framework bases all parameter estimations on peers identified within the unprotected group.
-This group typically possesses ample data points, effectively mitigating issues related to data scarcity and group imbalance, making our framework robust theoretically.
-We investigate the robustness of our peer-induced fairness framework by evaluating the percentages of unfairly treated ($PUT$) protected individuals or organisations and the invariant outcome ratio ($IOR$) under varying levels of imbalance.
-The imbalance ratio, $\omega$, is defined as the proportion of samples in the protected class:
-$\omega=\frac{\#(S=s_{-})}{\#(S=s_{+})+\#(S=s_{-})},$
-where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality of a set.
-A perfectly balanced dataset corresponds to $\omega=50\%$.
-The $PUT$ is calculated as the number of unfairly treated individuals or organisations divided by the total number of selected subjects in the experiments with different $\omega$.
-The $IOR$ is computed as the number of selected individuals or organisations in the experiment with $\omega$ that have unchanged predictive outcomes compared to the original experiment ($\omega=41.33\%$) divided by the number of commonly selected subjects in both the experiment with $\omega$ and the original experiment.
-In SMEs experiment, building upon the default settings outlined in Section 5, we explore the influence of varying imbalance ratios by randomly selecting subsets of the original dataset with controlled imbalance ratios.
-Specifically, we evaluate performance at imbalance ratios of $\omega=\{36.33\%,31.33\%,26.33\%,21.33\%,16.33\%,11.33\%\}$.
-By decreasing the percentage of micro-firms in these subsets, we assess the framework’s performance across different levels of imbalance.
-To mitigate the effects of randomness inherent in subset selection, the process is repeated five times.
-The detailed procedure is presented in Supplementary Materials.
-The results are visualised in Fig.
-5 and demonstrate the robustness of our framework.
-From the view of $PUT$, the small error bars across all the imbalance levels suggest the results across the five repetitions are highly consistent.
-This observation underscores the robustness of our “peer-induced framework” to imbalanced datasets.
-From the view of $IOR$, it is approximately 95% and remains stable across different imbalance levels.
-This aligns with our expectations, as the framework does not rely on data from the minority group but rather leverages information from the unprotected group, leading to inherent robustness.
-The small error bars suggest that for imbalance ratios greater than or equal to 16.33%, the results regarding $IOR$ are also highly consistent.
-Figure 5: Percentage of unfairly treated micro-firms and invariant outcome ratio at different group imbalance levels.
-The imbalance level is represented on the x-axis as a percentage, ranging from 11.33% to 36.33%.
-The left y-axis shows the percentage of unfairly treated micro-firms (blue line), while the right y-axis displays the invariant outcome ratio (red line) as the imbalance level changes from the initial level to other levels.
-These findings underscore the stability of our “peer-induced fairness” framework, distinguishing itself from others by effectively addressing the data scarcity and imbalance issues.
-Given the widespread nature of these issues, our framework holds considerable significance for researchers investigating algorithmic fairness and data imbalance.
-An alternative computation method is also provided in Supplementary Materials to ensure robustness.
-6.3 Explainable fairness discovery
-In our previous experiments, we were able to distinctly classify individuals from the protected group into two categories: fairly-treated and unfairly-treated groups.
-Our analysis now turns to those who were rejected while still fairly treated, to understand the reasons behind their rejections by comparing their features with those of their peers.
-Additionally, the “peer-induced fairness” framework allows us to provide a clear watch-out list of a series of features.
-Given the existence of accepted peers as the counterfactual instances with positive accessing finance outcomes, the micro-firm which is fairly treated should originally have the same outcomes.
-Our framework identifies the feature differences between each rejected while fairly treated micro-firm and its accepted peers by hypothesis testing, with details presented in Supplementary Materials.
-For each feature, we summarise the percentage of these micro-firms that perform significantly worse than their accepted peers.
-We consider some actionable and key features to identify and understand these discrepancies, as in Fig.
-6.
-The descriptions for each feature value are shown in Supplementary Materials.
-Results show that even though none of them have been rejected previously and only 25% of them perform worse on financial qualifications and written plans, banks generally prioritise the financial and business health of firms.
-75% of these micro-firms invest excessively in business innovation and have lower risk ratings.
-Besides, half of them invest in product/service development and have lower profits.
-The uncertain returns and high risks associated with innovation lead to the failure or commercial non-viability of most innovative products (Coad and Rao,, 2008; Hall,, 2002; Freel,, 2007), exacerbating already poor-performing risk indicators.
-The worse performance on these key features makes banks cautious about the long-term financial sustainability of these firms.
-It also reflects the capability of these micro-firms, negatively affecting their loan approvals.
-Figure 6: Comparative analysis of key features for rejected while fairly treated micro-firms vs.
-accepted peers.
-The x-axis represents the selected key attributes being analysed, including finance qualification (FQ), written plan (WP), previous turn-down (PT), loss or profit (LP), risk (RI), product/service development (PS), business innovation (BI).
-The y-axis represents the percentage of those micro-firms with significantly worse performance than their accepted peers on each feature.
-This exploration identifies the differences between micro-firms and their peers for each feature and summarises the percentage of micro-firms that perform worse on each feature.
-This explainable analysis not only enhances the transparency of our framework but also supports regulators and stakeholders in understanding the specific challenges most incapable micro-firms face and highlights the features that they need to watch out for and pay extra attention to.
-7 Conclusion
-In this paper, we introduce a novel fairness framework within a causal framework, termed “peer-induced fairness”, as a bias auditing tool for internal and external assessment, in a plug-and-play fashion.
-It applies the principles of counterfactual fairness, stipulating that the average treatment of an individual should align with that of their peers.
-We identify peers based on similar joint distributions but resort to $IC$ due to the unidentifiability of the counterfactual distribution.
-The framework requires equal treatment with peers.
-This approach effectively tackles data scarcity and group imbalance by utilising robust counterfactual statistics derived from well-represented peer groups, thereby ensuring more stable bias auditing.
-Besides, based on the essence of peer comparison, we could also provide an explainable watch-out list for those who receive unfavourable decisions due to insufficient capabilities, promoting transparency of our method.
-We have applied this framework to SMEs, but it has the potential for a generalisation to other research domains.
-By experimenting on SMEs data, our research findings reveal that by comparing micro-firms with their peers, banks and regulators can effectively audit algorithmic bias.
-Specifically, only 2.48% of micro-firms are treated fairly.
-41.51% and 56.40% of micro-firms are either discriminated against or privileged respectively.
-Even though some micro-firms benefit from algorithmic favouritism, it is essential to ensure equitable treatment across all applicants.
-Nearly half of the micro-firms experiencing extreme discrimination are rejected, with a rejection rate, compared to only 9.97% among their peers.
-This difference diminishes and becomes negative as discrimination lessens and shifts towards privilege.
-Up to 95% of micro-firms maintained consistent auditing results despite changing imbalance levels, demonstrating the stability of our framework with data scarcity and imbalance issues.
-Additionally, the approach highlights the key features that financial institutions need to pay more attention to and rejected micro-firms may need to address, whilst clearly fairly treating micro-firms.
-The comparison could distinguish the bias and incapability faced by micro-firms, helping banks and regulators understand the specific issues these firms encounter.
-Our research is significant for researchers who aim to scientifically audit algorithmic bias.
-This tool could perform with common data quality issues, like data scarcity and imbalance, ensuring accuracy and stability in measuring unfairness.
-Besides, our framework could also distinguish those incapable protected individuals from being biased individuals, preventing improving the treatment of less capable individuals at the expense of the treatment of capable, unprotected individuals.
-Besides, our empirical analysis is based on SMEs.
-Unlike previous studies that mainly focused on individual loans, our research extends the focus to the firm level.
-The protected feature is the firms’ size, differing from the traditional focus on person-level characteristics.
-This approach broadens the perspective of fairness.
-Despite the focus on SMEs loan approval, the fairness audit framework proposed can be applicable to other domains where algorithmic unfairness may occur, especially those suffering from group imbalance and data scarcity.
-In the domain of fairness research, class imbalance is also a crucial issue involving poor data quality.
-It refers to the imbalance on the target label, leading the model to favour the majority class, thereby affecting the overall performance of the model and the fairness measure.
-Previous studies discussing the impact of class imbalance on fairness focus on the education domain (Sha et al.,, 2022, 2023), with the issue remaining unexplored in the credit scoring domain.
-This is particularly important because different datasets exhibit significant variations in features, labels, missing values, and sample sizes.
-Besides, Iosifidis et al., (2021, 2019) have proposed fair models for addressing class imbalance, but there is still a lack of a fairness framework that explicitly considers class imbalance naturally.
-Exploring the impact of class imbalance on algorithmic fairness measures, and developing fairness criteria related to class imbalance, are crucial for all domains reliant on precise data-driven decision-making.
-Acknowledgments
-The authors of this manuscript would like to thank Prof.Raffaella Calabrese and Dr.Yizhe Dong for their assistance and support in the discussion and research direction.